A Complete Bibliography of Journal of the Royal Statistical Society, Series D (The Statistician):

2000–2009

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

13 October 2017 Version 1.01

Title word cross-reference

χ^{2} [291]. F [263]. κ [225]. N [387]. P [128]. t [263, 262]. t_{2} [474].

-Coefficient [225]. -Distribution [474]. -Test [262]. -Value [128].

100 [55]. 120 [150].

2001 [228].

3 [83].

Abdi [46]. Abramson [364, 364]. Accounts [107]. Accuracy [388]. Actuarial [42]. Acyclic [187]. Adaptive [279]. Additional [38]. Additive [3, 92]. Adèr [100]. Adjusting [430]. Adjustment [156, 507]. Advanced [237, 521, 217, 402, 420]. Advances [109]. Advent [157]. Age [32]. Ageel [277]. Ages [150]. Aggregates [190]. Agreement [153, 282]. Aiken [523]. Akers [495]. Alan [98]. Algebra [233, 487, 118]. Allgemeineren [461]. Allison [365]. Allocation [38]. Allometric [283]. Altman [14, 101]. Altmann [183]. Always [508]. America [71]. AMOS [330]. Analyses [60, 176]. Analysing [282]. Analysis [1, 277, 537, 94, 235, 172, 385, 320, 199, 286, 184, 110, 438, 366, 524, 473, 220, 416, 488, 523, 289, 278, 318, 67, 345, 66, 217, 87, 397, 134, 58, 68, 61, 197, 221, 309, 203, 138, 180, 211, 119, 483, 352, 205, 276, 204, 529, 74, 89, 297, 484, 488, 491, 137, 439, 50, 314, 145, 442, 425, 525]. Analyzing [107]. ANCOVA [380]. Anderson [71, 18]. Andorka [102]. Andreev [32]. Angle [262]. Annotated [301]. ANOVA [380, 136]. Anscombe [514]. Application [340, 506, 423]. Applications [1, 81, 148, 199, 452, 351, 533, 450, 330, 368, 221, 50, 470, 304, 241, 240, 418, 521, 540, 143, 171, 307, 416, 278, 236, 244, 166, 242, 338]. Applied [204, 315, 146, 541, 448, 447, 529, 211, 523, 305, 497, 105, 17]. Applying [243]. Approach [94, 204, 345, 283, 471, 124, 505, 226, 380, 88, 348, 218, 353, 379, 114, 238, 18, 451]. Approaches [146, 423, 439]. Approximating [371]. Approximation [370, 454]. Archaeology [174]. Areas [470]. Argyrous [232]. Arithmetic [51, 77]. Arithmetical [481]. Armitage [15]. Armstrong [401]. Arnold [131, 398]. Art [417]. Aspects [76, 171]. Assessing [227]. Assessment [56, 203, 226]. Assisted [20, 199]. Associated [19]. Association [188, 512, 289]. Assumption [65]. Asymmetric [507]. Asymptotic [499, 31]. Asymptotics [478, 249, 310]. Atkinson [297]. Attendance [290]. Auditing [355]. Auditors [471]. Autoregression [7]. Autoregressive [93]. Aven [47]. Averaging [124]. Ayeni [113].

B [483, 81, 387, 448, 462, 141, 313, 414, 327, 315, 378, 83, 421, 233, 529, 372, 145, 85, 405, 60, 140, 351, 541, 406, 46, 179, 108, 53, 217, 118, 134, 61, 52, 479, 330, 316, 176, 528, 131, 314, 201]. Babbie [24]. Back [34, 62, 90, 121, 151, 186, 223, 253, 265, 296, 328, 363, 400, 437, 476, 520, 156]. Backofen [200]. Baines [398]. Balakhrishnan [143]. Balance [69]. Band [156, 507]. Banks [83]. Bapat [233]. Barden [82]. Bardi [132].

Barnard [435]. Barnett [16, 477]. Bartlett [325]. Based [39, 122, 259, 521, 367, 7, 241]. Basic [310, 217, 330, 538, 314]. Basics [308, 490, 413]. Batched [226]. Bates [210]. Bather [298]. Batting [8]. Baumert [146]. Bayes [356, 355, 4]. Bayesian [422, 369, 379, 22, 451, 201, 94, 42, 64, 506, 427, 471, 505, 266, 110, 412]. Be [36, 245]. Beat [282]. Bedford [438]. Beginners [237]. Beginning [335]. Behavioral [523, 377, 492]. Behavioural [100, 432]. Behling [163]. Beltrami [133]. Bendat [197]. Benjamin [515]. Benson [329]. Beran [72]. Bernard [515]. Bernardo [266]. Bernstein [366]. Berry [348]. Bertollini [56]. Best [236, 370]. Bet [96]. Bets [382]. Better [508]. Betting [284, 96]. between [285]. Bias [319]. Biased [472]. Biggeri [56]. Billingsley [73]. Bilodeau [103]. Binary [509, 93, 506]. Binder [344]. Binomial [127]. Biological [424]. Biologist [469]. Biology [200, 139]. Biometrika [459]. Biopharmaceutical [135]. Biostatistics [15, 452, 489]. Bisgaard [500]. Bivariate [473, 393]. Bland [367, 234]. Blischke [198]. Bloomfield [235]. Blossfeld [439]. Blue [465]. Bobko [368]. Bogaert [521]. Böhning [56, 199]. Book [109, 250, 57, 119, 483, 72, 16, 26, 79, 81, 248, 268, 277, 380, 387, 448, 237, 200, 299, 308, 352, 537, 304, 344, 365, 462, 174, 443, 482, 536, 340, 149, 235, 447, 56, 219, 407, 534, 15, 47, 88, 141, 172, 266, 313, 382, 415, 246, 116, 385, 148, 205, 14, 20, 199, 241, 25, 298, 384, 422, 461, 73, 198, 414, 493, 271, 184, 135, 48, 113, 170, 243, 276, 460, 267, 27, 107, 167, 204, 311, 337, 342, 408, 110, 315, 242, 378, 412, 147, 30, 80, 83, 215, 240, 115, 421, 233]. Book [55, 371, 529, 76, 273, 372, 438, 459, 366, 74, 144, 54, 24, 89, 335, 51, 71, 145, 146, 207, 208, 369, 418, 521, 85, 341, 209, 234, 405, 452, 524, 487, 540, 530, 539, 60, 103, 332, 336, 348, 214, 297, 78, 84, 106, 133, 140, 142, 143, 165, 169, 171, 183, 185, 211, 218, 220, 239, 245, 274, 275, 301, 302, 307, 329, 343, 347, 351, 353, 367, 373, 379, 402, 535, 541, 403, 404, 406, 411, 416, 417, 423, 441, 444, 445, 455, 456, 458, 477, 481, 484, 488, 491, 498, 523, 526, 527, 531, 532, 533, 206, 306, 168]. Book [164, 270, 339, 46, 173, 278, 376, 249, 310, 381, 454, 494, 236, 349, 244, 374, 100, 86, 150, 179, 108, 114, 19, 450, 377, 175, 53, 177, 29, 136, 232, 305, 410, 446, 497, 82, 345, 346, 370, 420, 105, 312, 495, 350, 383, 178, 216, 217, 87, 238, 117, 17, 489, 59, 213, 375, 22, 118, 442, 496, 522, 23, 111, 134, 58, 364, 18, 499, 61, 52, 139, 202, 212, 479, 77, 330, 334, 388, 424, 485, 490, 31, 182, 492, 457, 272, 338, 451, 101, 538, 279, 386, 132, 197]. Book [247, 251, 453, 478, 480, 102, 137, 163, 368, 419, 449, 486, 28, 104, 303, 316, 181, 331, 413, 440, 333, 409, 425, 542, 439, 501, 33, 49, 221, 309, 176, 203, 525, 32, 75, 166, 401, 528, 112, 210, 21, 138, 180, 500, 269, 131, 50, 314, 201, 300]. Boots [351]. Bootstrap [357, 5, 104]. Bosker [217]. Box [500]. Bradley [48]. Brass [360]. Breadline [339]. Brenner [103]. Brewer [479]. Brian [11]. Briggs [236]. Brown [17]. Browne [312]. BSE [76]. Building [510, 182]. Bürger [267]. Burnham [18]. Busines [193]. Business [342, 259, 193, 464, 208, 275, 419, 303]. Byrne [330].

C [81, 299, 352, 174, 407, 534, 88, 172, 313, 384, 422, 493, 184, 135, 48, 113, 327, 83, 115, 421, 76, 74, 209, 524, 539, 84, 106, 211, 245, 302, 329, 411, 416, 445, 206, 164, 236, 450, 177, 305, 446, 17, 22, 58, 210, 138, 500, 131]. C.-Y [113]. Cabrera [440]. Calculated [382]. Calculations [2]. Calculus [540, 118]. Calibration [392]. Cam [310]. Cambridge [406, 468]. Campbell [498]. Can [245]. Canonical [50]. Carlo [344, 115, 371, 527]. Carne [82]. Case [447, 290, 228, 387]. Case-Study [228]. Casella [115]. Cases [227]. Casino [288]. Castillo [131]. Categorical [205, 58]. Causal [439, 188].

Causation [384]. Cavazos [106]. Cavazos-Gaither [106]. Census [71, 228]. Census-Taking [71]. Centenary [19]. Centiles [3]. Centuries [411]. Chain [293, 292]. Chance [133, 172, 141]. Chances [274]. Changepoint [126]. Chaos [341]. Chapernowne [516]. Characteristic [430]. Characteristics [512]. Characterizations [474]. Charts [507]. Chatfield [74, 299]. Chatterjee [134]. Checking [471]. Chemical [180]. Chen [478]. Chernick [104]. Chiles [49]. Choice [422]. Chow [135]. Christakos [521]. Christensen [402]. Christopher [326]. Chui [351]. Chung [441]. Classification [431, 79]. Climate [385, 87]. Clinical [81, 482, 36, 191, 496, 522, 405]. Clive [409]. Clote [200]. Cluster [358, 483]. Cobweb [6]. Codes [261]. Coefficient [263, 225, 129]. Coefficients [322, 122]. Coggan [522]. Cognition [86]. Cohen [523]. Coles [307]. Collected [409]. Collection [137, 314, 20]. Collett [75]. Collinear [318]. Collins [74]. Colton [15]. Combinatorial [332]. Combinatorics [417]. Committees [482]. Communication [331]. Communications [533]. Comparative [16]. Comparison [65, 256, 285, 191]. Comparisons [96]. Competition [69]. Complex [261]. Complicated [510]. Component [488]. Computation [263, 423]. Computational [200]. Computer [341, 261, 138, 366, 20, 199]. Computer-Assisted [199]. Computerized [279]. Computers [382]. Computing [105]. Concepts [351, 416, 377, 495, 330, 492, 310]. Conceptual [432]. Conditional [122, 131]. Conference [193]. Confessions [255]. Confidence [357]. Congdon [369]. Conniffe [19]. Conquering [460]. Consensus [35]. Consistent [393]. Constrained [14]. Constraints [38]. Construction [3]. Consulting [331, 440]. Contact [319]. Contemporary [71].

Contingency [66, 188, 6]. Continuous [190]. Contour [32]. Control [156, 500]. Controversy [275, 35, 303]. Conventional [393]. Convergence [73]. Cook [105]. Cooke [438]. Corbelled [64]. Cormier [493]. Correction [250]. Correlated [509]. Correlation [263, 322, 472, 523, 129, 243, 336, 368]. Correspondence [324, 9, 434, 159]. Corrigendum [222]. Cortina [136]. Costs [507]. Council [97, 230, 359, 513]. Count [220]. Counting [417]. Counts [71]. Couper [20]. Coupling [219]. Course [352, 172, 205, 414, 441, 212, 492, 528]. Courtroom [273]. Covariance [510]. Covariates [430]. Cowles [403]. Cox [248, 268, 300, 459]. Coxon [137]. Crano [479]. CRC [185]. Cricket [293, 8, 287]. Criteria [395]. Criterion [96]. Cross [192, 496]. Cross-over [496]. Cross-Sectional [192]. Crunch [460]. Cumulative [39, 428]. Current [470]. Curricula [193]. Curve [166]. Curves [430]. Cuthbert [10]. Cvitani [343]. Cyganowski [480].

D [109, 79, 268, 352, 537, 344, 365, 462, 482, 56, 415, 199, 25, 198, 493, 276, 27, 408, 315, 80, 83, 459, 51, 146, 487, 540, 530, 103, 336, 185, 211, 239, 245, 275, 329, 347, 373, 404, 455, 458, 498, 532, 339, 46, 278, 376, 236, 244, 86, 114, 19, 450, 53, 305, 82, 420, 105, 312, 216, 213, 522, 23, 18, 479, 77, 101, 386, 478, 449, 486, 303, 316, 33, 203, 75, 210, 500, 269, 201, 300]. Daley [404]. Daniel [138, 10]. Daniels [160]. Data [119, 127, 248, 153, 509, 172, 227, 205, 286, 48, 122, 292, 342, 366, 89, 146, 473, 211, 129, 397, 191, 364, 61, 139, 92, 197, 137, 32, 75, 138, 391, 314, 365, 220, 442, 58, 364, 33, 525, 138, 482, 416]. Davey [239]. Davey-Smith [239]. David [516, 301, 519]. Davis [524]. Dead [156, 507]. Decathlon [286]. Deciding [152]. Decision [14, 408, 412, 211, 373, 451, 124, 298]. Decisions [298]. Default [356]. Delayed [507]. Delfiner [49]. Demand [290]. DeMets [482]. Demographic [32]. Density [332]. Dependence [336]. Dependent [322]. Derek [518]. Derivatives [306, 271]. Derr [331]. Design [222, 119, 268, 184, 204, 89, 125, 111, 500, 300, 180]. Designed [509]. Designs [387, 123, 136]. Desrosières [164]. Detecting [428]. Detection [256, 191, 189]. Determination [187, 129, 427]. Determining [358, 226]. Deterministic [371, 507]. Deutsch [370]. Developments [464]. Devroye [332]. Dey [21, 201]. Dhrymes [333]. Diagnostic [227, 297]. Diagnostics [127]. Diagrams [6, 351]. Diamantopoulos [334]. Diamond [335]. Dictionary [106, 406, 52]. Did [282]. Difference [224, 427]. Differential [242, 30, 132, 480]. Diggle [442, 525]. Digits [95]. Dijk [530]. Dillman [486]. Dimensional [80]. Direct [270]. Directed [187]. Directional [147]. Directions [428]. Disclosure [259]. Discordancy [395]. Discovering [237]. Discovery [500]. Discrete [183, 156, 247, 33, 484]. Discrete-Event [484]. Disease [199, 431, 56]. Dispersion [3, 92]. Distance [348]. Distribution [263, 258, 533, 168, 474, 260]. Distributional [171]. Distributions [307, 140, 183]. Diversities [503]. DNA [293]. Dobson [526]. Domes [64]. Done [245]. Donnelly [76]. Dorfman [353]. Dorling [51, 239, 77]. Dose [222, 125]. Double [471]. Double-Checking [471]. Doucet [527]. Dr [195]. Draper [312]. Drawing [181]. Drift [507]. Drouet [336]. Due [472]. Dugard [387]. Duley [405]. Dunn [529]. Durrett [165]. Dutch [69]. Dynamic [298, 284, 44]. Dynamical [148, 148]. Dynamics [507, 537].

Econometric [540, 84, 304, 220]. Econometrics [241, 409, 242, 333]. Econometrics-Collected [409]. Economic [167, 275, 303]. Economics [533, 173]. Economists [142]. Economy [320]. Edelman [46]. Edge [256]. Editorial [390, 281]. Education [377, 492]. Edward [295]. Edwards [301, 481, 269]. Edwin [161]. Effect [136]. Effective [193, 331]. Effects [156, 189, 210]. Efficient [396]. Efromovich [166]. Elementare [461]. Elementary [480]. Elements [302, 28]. Eliciting [261]. Ellenberg [482]. Elliott [236]. Eltinge [486]. Embretson [202]. Emmett [11]. Empirical [530, 338]. Encounters [172]. Encyclopedia [15, 135, 83]. Engineering [533, 450, 216, 314]. Engineers [53]. English [284, 290, 504]. Ennos [139]. Entropy [22]. Enumerative [417]. Environmental [415, 378, 424, 349]. Epidemic [404]. Epidemics [76]. Epidemiological [192, 364]. Epidemiology [236, 119, 89]. Epiphanies [329]. Equality [322]. Equation [376, 212, 330]. Equations [480, 138, 30]. Erickson [22]. Ernest [195]. Erritt [475]. Error [431]. Errors [67, 156]. Essays [409]. Essential [342, 53, 213]. Essentials [165]. Estimating [225, 67]. Estimation [222, 190, 126, 125, 427, 41, 394, 5, 166, 260, 332, 59]. Europe [339]. European [320, 289]. Evaluation [208, 512]. Evander [282]. Evans [337, 371, 140]. Event [439, 484]. Everitt [483, 529, 52, 141, 60, 406, 176, 528]. Everson [488]. Evidence [116, 367, 319, 239]. Evidence-Based [367]. Evolution [193]. Evolutionary [469]. Exact [263, 357, 427]. Example [134]. Examples [146]. EXCEL [257]. Exceptional [108]. Exercises [487]. Expectation [251]. Expectations [337]. Experiment [508]. Experimental [387, 91, 123, 111]. Experimenting [246].

Experiments [184, 268, 300]. Expert [91]. Exploratory [211, 289]. Exploring [415, 24]. Exponential [154, 511]. Extending [248]. Extension [260]. Extensions [376]. Extreme [307, 307].

F

[56, 172, 266, 385, 184, 378, 412, 24, 452, 211, 301, 481, 179, 305, 410, 370, 87, 118, 32, 138]. Face [80]. Factor [66, 358]. Factorial [21]. Failure [2]. Fairly [468]. Farrell [405]. Fayers [203]. Feed [156]. Feed-Back [156]. Ferguson [76]. Field [521, 420, 237]. Field-Based [521]. Fields [30]. Fienberg [71]. Finance [149, 530, 533, 29]. Financial [320, 271, 84, 425, 542, 306]. Finite [394, 247, 347, 353]. Finkelstein [372]. First [172, 414, 292, 212, 453, 462]. First-Order [292]. First-Passage [453]. Fischer [270]. Fisher [468, 469, 466, 467, 465]. Fit [323]. Fitting [293, 292, 321, 138]. Fitzgerald [407]. Fixed [277]. Flannery [421]. Fleming [482]. Flows [190]. Flshman [484]. Fluctuations [311]. Folded [511]. Football [284, 43, 512]. Force [150]. Forecast [419]. Forecasting [221, 401, 299]. Forecasts [7]. Foreign [270]. Foreword [465]. Formulae [185]. Forward [127]. Foundations [438, 376]. Fouque [271]. Fourier [235]. Fox [272]. Fractional [21]. Frank [514]. Franses [530]. Frederick [518]. Freitas [527]. French [408, 373]. Front [13, 45, 70, 99, 130, 162, 196, 231, 254, 280, 317, 354, 426, 463, 502, 544]. FRS [325]. Frumkin [167]. Fuente [142]. Function [348]. Functions [263, 521]. Fundamentals [388, 478].

G [352, 537, 172, 246, 271, 27, 204, 337, 80, 115, 55, 529, 144, 521, 341, 332, 183, 239, 347, 417, 423, 456, 484, 523, 532, 310, 236, 100, 86, 450, 377, 232, 82, 312, 350, 213, 375, 22, 23, 212, 492, 451, 101, 197, 102, 439, 500]. Gaither [106, 106]. Gambling [382].

Gamerman [114]. Games [132]. Gani [404]. Garson [23]. Gastwirth [273]. Gaussian [261]. Gawen [516]. Geary [19]. Geer [338]. Generalizations [533]. Generalized [27, 526, 272, 450, 446, 238, 59, 201]. Generation [275, 303]. Genetic [491]. Geneticist [468]. Genetics [468]. Geographic [418]. Geohydrology [17]. Geometrical [171]. Geometry [242]. George [295, 435]. Geostatistical [534]. Geostatistics [49]. Geostatistics-Modeling [49]. Gerber [61]. Ghosh [201]. Ghysels [409]. Giblin [80]. Gifts [503]. Gill [238]. GIS [521]. Glance [32, 209]. Glaz [143, 531]. Gliner [204]. GLM [380]. Gloeckner [350]. Glymour [384]. Goldstein [312]. Goldthorpe [374]. Golf [289]. Good [78]. Goodness [323]. Goodness-of-Fit [323]. Gordon [80, 239, 339, 79, 245, 527]. Gore [264]. Gorte [179]. Grafen [485]. Grambsch [248]. Granger [409]. Graphical [54, 226, 269]. Graphics [105, 61]. Graphs [187, 412]. Graubard [145]. Greenfield [443]. Griego [350]. Group [81]. Groves [486]. Grundbegriffe [461]. Guide [387, 344, 141, 232, 312, 23, 334, 92, 453, 419, 331, 340, 104, 167].

H [26, 277, 219, 385, 148, 398, 460, 311, 421, 366, 418, 530, 539, 211, 301, 353, 46, 173, 278, 374, 100, 114, 450, 29, 136, 82, 312, 87, 364, 182, 279, 386, 478, 181, 413, 439, 501, 50]. H.-P [439]. Hadi [134]. Hahn [53]. Haigh [274]. Hails [485]. Halley [24]. Hallinan [80]. Hamada [184]. Handbook [537, 60, 455, 118, 176, 401]. Handcock [168]. Handling [139]. Härdle [340]. Harold [362]. Harrell [410]. Harville [487]. Hassenzahl [415]. Hastings [140]. Hate [214]. Heagerty [442, 525]. Healey [24]. Health [415, 113, 145, 239, 232, 56]. Healy [312, 444]. Heathcote [302]. Henry [361, 160]. Herbert [12]. Herrmann [86]. Hesketh [60, 176]. Heterogeneous [472]. Heteroscedasticity [94]. Heyde [411]. Hidden [257]. Hierarchical [355]. Hill [500]. Hirschberg [275, 303]. Historical [403]. History [301, 164, 439]. Hlávka [340]. Hoaglin [211]. Hodson [107]. Hollander [25]. Holly [304]. Holyfield [282]. Honkapohja [337]. Hoover [341]. Hosmer [305]. Hoyle [26]. Hsiao [445]. Hsu [110]. Hundred [459]. Hungary [102]. Hunt [306]. Hurd [342]. Hutcheson [27]. Hyvärinen [488].

Ian [436]. Idea [481]. II [346]. Images [256]. Improved [192, 323]. Improvement [178]. Improving [388]. Including [105]. Incomplete [397]. Incorporating [226]. Independence [393]. Independent [488, 488]. Indexed [144]. Indicators [167]. Indices [430]. Indrayan [489]. Industrial [270, 368]. Industry [447, 91, 506]. Inequalities [245, 239]. Inference [192, 241, 37, 126, 353, 458, 494, 114, 18, 499, 501, 172, 16, 414, 85]. Inferences [153, 392, 181]. Inferential [128]. Inflation [358]. Influence [227]. Influential [189]. Informal [141]. Information [120, 252, 389, 543, 313, 20, 356, 418, 18]. Information-Theoretic [18]. Initial [93]. Inner [370]. Instruction [364]. Instruments [163]. Insua [408, 373]. Insurance [29]. Integrals [371]. Integrated [204, 114]. Integration [374]. Interactive [356]. Interconnected [293]. Interdisciplinary [110]. Interest [343]. Interface [109]. Intermediate [528, 117]. Intermittent [313]. International [224]. Interobserver [153]. Interpretation [350, 364, 203]. Interpretations [169]. Interrater [282]. Interval [128, 357]. Intervallwahrscheinlichkeit [461]. Intervals [40, 7]. Intraclass [322]. Introducing [380, 334]. Introduction [149, 298, 335, 171, 307, 526, 527, 350, 217, 118, 235, 404, 111, 215, 74, 234, 457, 221, 269]. Introductory [27, 23, 173]. Invariant [345]. Investment [270]. Irish [19]. Issues [24, 146, 432]. Item [457, 202]. Ivanoff [144].

J [109, 57, 72, 387, 448, 536, 56, 407, 172, 266, 313, 241, 298, 271, 184, 398, 48, 243, 460, 204, 342, 110, 378, 273, 74, 24, 335, 71, 146, 418, 348, 214, 143, 211, 218, 274, 343, 347, 367, 379, 404, 444, 445, 488, 498, 523, 531, 532, 533, 161, 306, 454, 236, 374, 100, 86, 150, 108, 175, 29, 136, 497, 345, 312, 217, 238, 117, 22, 442, 58, 364, 334, 424, 272, 538, 197, 480, 419, 486, 331, 440, 333, 409, 542, 49, 309, 525, 32, 401, 112, 210, 500, 131, 50, 314]. J.-F [56]. J.-P [271]. Janacek [375]. Jefferies [335]. Jennison [81]. Jensen [47, 412]. Jeune [108]. Job [275, 303]. Job-Generation [275, 303]. Jobe [314]. John [475, 294, 532, 19, 295]. Johnson [532]. Jordon [54]. Jouini [343]. Journal [390, 229]. Jr. [410]. Jupp [147].

K.-Y [442, 525]. Kahane [413]. Kakizawa [499]. Kale [414]. Kammen [415]. Kanji [55]. Kannisto [150]. Kaplan [376]. Karhunen [488]. Kay [109]. Kelly [96]. Kempthorne [194]. Kempton [517]. Kendall [82]. Kennedy [306]. Keuzenkamp [501]. Key [313]. Khrennikov [169]. Klinke [340]. Kloeden [480]. Knight [170]. Knockout [285]. Knowledge [226]. Known [468]. Kokoska [185]. Kolosi [102]. Konzept [461]. Korn [145]. Kotz [83, 307, 533, 336]. Kozubowski [533]. Krause [308, 490, 447]. Krzanowski [309].

L [482, 407, 88, 148, 493, 311, 80, 83, 273, 145, 418, 85, 405, 332, 535, 441, 445, 523, 310, 175, 59, 442, 486, 28, 413, 525, 180]. Laboratory [191]. Lancaster [361]. Landau [483, 344]. Lange [491]. Langford [312, 436]. Lantuejoul [534]. Laplace [533]. Large [510, 36, 164, 28]. Large-Sample [28]. Lauritzen [535]. Law [163]. Lawson [56]. Lawyers [372]. League [284, 68]. Learning [193, 337, 54]. Lecture [19]. Leese [483]. Lehmann [28]. Lele [345]. Lemeshow [305]. Lennox [282]. Leonard [110, 205]. Lesaffre [56]. Lesson [113]. Lethal [222, 125]. Levin [372, 111]. Lewis [282, 312]. Liang [442, 525]. Lieshout [240]. Life [466, 100, 432, 485, 203, 133]. Lifetime [4]. Likelihood [116, 494, 41, 206, 494, 381]. Limitation [259]. Limited [8, 287]. Lindsey [57, 536, 112]. Linear [429, 27, 233, 402, 526, 450, 446, 238, 321, 118, 394, 201, 233, 530, 177]. Link [433]. Liptser [346]. LISREL [334]. Little [146, 486]. Lloyd [58]. Loader [206]. Local [206]. Location [5]. Logistic [305, 123, 305]. Lomax [377, 492]. London [468]. Long [72, 152]. Long-Memory [72]. Longevity [108]. Longitudinal [146, 442, 189, 525]. Look [154]. Lottery [291]. Lugosi [332]. Lunneborg [416]. Lwanga [113].

M [109, 119, 483, 248, 277, 308, 304, 149, 266, 415, 20, 241, 25, 184, 243, 242, 240, 325, 371, 76, 372, 459, 54, 89, 71, 208, 521, 234, 103, 297, 140, 171, 239, 343, 353, 367, 379, 403, 404, 423, 444, 498, 168, 86, 136, 420, 312, 499, 479, 330, 490, 132, 247, 137, 486, 104, 409, 501, 221, 203, 210, 500, 131, 314]. Machin [203]. Mackey [475]. Macroeconomics [337]. Making [43, 224, 451, 193, 364]. Mallick [201]. Management [207, 342, 378, 343, 368]. Managers [532]. Mangan [342]. Manly [378]. Manual [364]. Many [505]. MAPLE(R) [480]. Mapping [56, 199]. Maps [32]. Marcoulides [212]. Mardia [147]. Mari [336]. Mariano [241]. Markets [271]. Markov [14, 293, 292, 512, 431, 257, 240]. Marriott [242]. Marsh [493]. Martin [417].

Martingales [144]. Matches [293, 284, 68]. Mathai [171]. MATHEMATICA [118, 41]. Mathematical [149, 382, 267, 302, 329, 481, 173, 454, 244, 182, 170, 142, 491]. Mathematics [133, 247, 333]. MATLAB [53]. Matrices [540, 444]. Matrix [487, 540]. Matter [34, 62, 90, 121, 13, 45, 70, 99, 151, 186, 223, 253, 130, 162, 196, 231, 265, 296, 328, 363, 254, 280, 317, 354, 400, 437, 476, 520, 426, 463, 502, 544]. Mátyás [59]. Maunder [518]. Maximum [41, 22]. McAleer [501]. McClave [172]. McCulloch [446]. McDougall [440]. McGee [221]. McLachlan [347]. Mean [420, 3, 394, 92]. Measurement [339, 197]. Measurements [524, 112]. Measures [129, 73]. Median [222, 125]. Medical [536, 234, 451, 101, 75, 209, 489]. Medicine [367]. Meer [179]. Mellenbergh [100]. Memory [72]. Merzbach [144]. Meta [199]. Meta-Analysis [199]. Metcalfe [207]. Method [259, 59]. Methodology [100, 432, 464]. Methods [81, 443, 241, 356, 276, 204, 110, 524, 332, 348, 142, 491, 168, 381, 236, 282, 178, 216, 191, 479, 388, 538, 104, 166, 25, 115, 371, 438, 78, 477, 527, 420, 22, 132]. Michael [475]. Microdata [259]. Mielke [348]. Migon [114]. Miles [243]. Milestones [504]. Millard [447, 349]. Miller [418]. Mills [84]. Missing [227, 365]. Misuses [48]. Mixed [277, 446, 189, 210]. Mixed-Effects [210]. Mixture [347, 391]. Mixtures [199]. MLwiN [312]. Model [94, 429, 187, 37, 292, 259, 512, 355, 4, 124, 391, 248, 18, 182]. Model-Based [259]. Modeler [218]. Modeling [382, 307, 484, 410, 330, 439, 402, 445, 217, 146, 451, 542, 49]. Modelling [304, 88, 198, 284, 64, 84, 404, 376, 494, 92, 448, 369, 212, 501, 269, 248, 75]. Models [293, 536, 47, 154, 122, 76, 431, 530, 473, 142, 257, 450, 177, 123, 238, 3, 538, 92, 189, 188, 210, 201, 277, 493, 48, 27, 233, 54, 347, 526, 446, 131, 33, 112]. Modern [388, 315, 541, 485]. Molecular [200]. Molenaar [457]. Moment [329]. Moments [155, 59, 50]. Monitoring [482]. Monte [344, 371, 527, 115]. Montgomery [450]. Morgan [350, 448, 204]. Morimune [445]. Morrell [419]. Morris [168]. Mort [449]. Mortality [150]. Most [39]. Mosteller [211]. Mouhammed [173]. Mukerjee [21]. Multifactor [138]. Multilevel [146, 67, 123, 217, 217]. Multilocation [355]. Multinomial [323]. Multiple [523, 272]. Multistate [431]. Multivariate [392, 320, 27, 529, 74, 103, 318, 129, 497, 17, 191, 428, 309, 452]. Multivariate-Univariate [392]. Multiway [6]. Mumford [80]. Muriel [264]. Murthy [198]. Musiela [343]. Mutation [267]. Myers [450]. Myth [275, 303].

N [268, 198, 27, 167, 315, 240, 76, 366, 214, 140, 143, 351, 527, 236, 349, 86, 346, 216, 316, 409, 300]. Nadarajah [307]. Naish [164]. Narratives [374]. National [319, 464]. Naus [531]. Neerchal [349]. Networks [109, 412, 23, 478, 215, 46]. Neural [109, 46, 23]. NLMIXED [321]. Noisy [256]. Nolan [244]. Nominal [153]. Non [2, 122, 395, 319, 321, 449, 486, 530]. Non-Contact [319]. Non-Linear [321, 530]. Non-Official [449]. Non-Randomized [2]. Non-Response [486]. Non-Spurious [395]. Non-Stationary [122]. Nonlinear [536, 407, 445]. Nonparametric [455, 428, 457, 272, 505, 25, 166]. Nonsensical [472]. Nonstationary [407]. Normal [129]. Normality [539]. Note [263, 429, 358]. Notes [466]. Nouri [136]. Novel [260]. Number [228]. Numbers [460, 164, 374]. Numerical [132, 421].

O [113, 372, 244, 350, 163]. Obituary [264, 475]. Observations [189]. Observed [190]. Off [507]. Off-Target [507]. Office [464]. Official [228, 449]. Oja [488]. Okabe [351]. Oliver [361]. Olsen [490]. Olson [308]. Olympics [283]. Ombach [480]. One [508, 459, 228, 540, 66, 498]. One-Number [228]. Operating [430]. Opinion [91]. Opper [420]. Optimal [288, 123]. Optimization [148, 198, 184, 478]. Optimum [38]. Option [343]. Options [149]. Order [292, 133]. Organizational [368]. Orton [174]. Oscar [194]. Other [149, 163]. Outcome [153, 290]. Outlier [191]. Outlines [113]. Overs [8, 287].

P [248, 387, 200, 344, 365, 235, 447, 407, 15, 148, 20, 384, 422, 73, 198, 271, 276, 242, 147, 80, 115, 421, 369, 521, 530, 348, 78, 523, 306, 270, 339, 236, 349, 383, 178, 216, 442, 111, 18, 202, 182, 251, 480, 137, 368, 333, 439, 49, 203, 525]. Pampel [305]. Panel [43]. Pantazis [245]. Papanicolaou [271]. Papers [409, 19]. Paradigm [116]. Parameter [184, 260]. Parameters [192]. Parametric [414, 455]. Pareto [4]. Parmigiani [451]. Parthasarathy [132]. Partial [30, 129]. Pascal [481]. Passage [453]. Patrick [11]. Patterns [80]. Pawitan [494]. Peacock [140, 367]. Pécseli [311]. Peel [347]. Pelosi [208]. Peña [352, 500]. People [214]. Percival [276]. Performance [395, 478]. Permutation [452, 348]. Personal [366]. Perspective [403, 102, 309, 201]. Pesarin [452]. Petrie [209]. Pharmaceutical [447, 35]. Phillips [175]. Philosophy [63]. Physical [311]. Physics [344]. Piersol [197]. Pinheiro [210]. Pioneer [535]. Planning [184]. Plans [21]. Plaque [465]. Plewis [312]. PLUS [308, 315, 490, 447, 349, 495, 542, 210]. Podgórski [533]. Point [240]. Poisson [65, 473, 397]. Poker [288]. Politics [71, 164, 51, 77]. Pollock [537]. Poly [37]. Poly-Weibull [37]. Pooling [472]. Pools [43]. Population [353, 394, 32]. Postgraduates [443]. Poverty [339].

Powell [445]. Power [323, 395, 388].

Powerful [39]. Practical

[387, 510, 146, 18, 246, 369, 375, 424, 101]. Practice [488, 306, 522]. Practitioner [104]. Practitioners [401]. Pragmatic [255]. Prakasa [85]. Precalculus [118]. Precedence [40]. Precursors [229]. Prediction [227, 40, 384, 198, 284, 353, 68, 4, 7]. Prehistoric [64]. Premium [42]. Presence [356]. Press [421, 379]. Price [134]. Pricing [343]. Pridmore [98]. Primer [305, 279]. Principles [42, 418, 477, 488, 57, 479, 309, 401]. Prior [356, 226]. Priors [261]. Probabilistic [438]. Probabilities [287]. Probability [141, 40, 73, 171, 183, 185, 441, 480, 50, 169, 274, 302, 456, 251]. Problem [415, 93, 126]. Problems [392, 282, 163]. Procedure [321, 428]. Procedures [192, 65, 455, 197]. Process [512, 507, 261, 338]. Processes [240, 29, 346, 72, 14, 165, 453]. Processing [537, 407]. Product [370, 394]. Products [226]. Professional [289]. Professor [361, 360, 516, 327, 468, 325, 515, 194, 294, 435]. Programming [298, 484, 383, 330, 182, 316]. Progression [431]. Project [228]. Pronzato [148]. Proof [329]. Proportions [427, 511]. Psychologists [528, 52, 202]. Psychology [403, 368]. Public [56].

Quality [178, 500, 203]. Quantifying [91]. Quantile [470]. Quantitative [107]. Questionnaires [163]. Questions [367]. Queueing [478]. Quotations [106].

R [26, 268, 200, 352, 482, 56, 407, 116, 241, 384, 198, 493, 271, 267, 107, 215, 233, 438, 459, 165, 211, 218, 220, 302, 353, 402, 444, 456, 488, 278, 454, 86, 150, 450, 377, 446, 345, 346, 105, 495, 217, 22, 18, 139, 388, 485, 492, 102, 486, 104, 409, 425, 221, 21, 180, 300]. R. [468, 466, 467]. Rabe [60, 176]. Rabe-Hesketh [60, 176]. Raghavan [132].

Raiffa [211]. Rain [287]. Random [277, 2, 95, 133, 285, 346, 4, 30, 197]. Randomization [387]. Randomized [2, 358]. Rao [85]. Rasbash [312]. Rates [343, 397]. Rating [44, 158]. Ratio [394]. Ratio-Product [394]. Ratios [155]. Raykov [212]. Read [83]. Readings [301]. Reasoning [164]. Receiver [430]. Recipes [421]. Recombination [267]. Recovery [270]. Redfern [161]. Redner [453]. Rees [213]. Reference [3]. Refusal [319]. regarding [393]. Regeneration [219]. Regression [127, 429, 93, 493, 243, 126, 297, 423, 523, 206, 305, 105, 394, 368, 189, 470, 305, 272, 410, 134, 413]. Regression/Correlation [523]. Reid [268, 300]. Reise [202]. Related [430, 17]. Relating [111]. Relationships [321]. Relative [168]. Reliability [216, 47, 88, 198]. Remark [396]. Remote [179]. Rencher [177]. Rental [92]. Repeated [524, 129, 112]. Report [97, 230, 359, 513]. Resampling [416, 78]. Research [374, 432, 232, 163, 470, 26, 385, 86, 350, 87, 496, 479, 101, 75, 180, 443, 204, 208, 100]. Researchers [110, 383, 401]. Response [66, 202, 457, 486]. Responses [318]. Responsible [224]. Restructuring [390]. Results [43]. Retrospective [68]. Revealing [57]. Reversibility [341]. Review [109, 57, 119, 483, 72, 16, 26, 79, 81, 248, 268, 277, 380, 387, 448, 237, 200, 299, 308, 352, 537, 304, 344, 365, 462, 174, 443, 482, 536, 340, 149, 235, 447, 56, 219, 407, 534, 15, 47, 88, 141, 172, 266, 313, 382, 415, 246, 116, 385, 148, 205, 14, 20, 199, 241, 25, 298, 384, 422, 461, 73, 198, 414, 493, 271, 184, 135, 48, 113, 170, 243, 276, 460, 267, 27, 107, 167, 204, 311, 337, 342, 408, 110, 315, 242, 378, 412, 147, 30, 80, 83, 215, 240, 115, 421, 233, 55]. Review [371, 529, 76, 273, 372, 438, 459, 366, 74, 144, 54, 24, 89, 335, 51, 71, 145, 146, 207, 208, 369, 418, 521, 85, 341, 209, 234, 405, 452, 524, 487, 540, 530, 539, 60, 103, 332, 336, 348, 214, 297, 78, 84, 106, 133, 140, 142, 143, 165, 169, 171, 183, 185, 211, 218, 220, 239, 245, 274, 275, 301, 302, 307, 329, 343, 347, 351, 353, 367, 373, 379, 402, 535, 541, 403, 404, 406, 411, 416, 417, 423, 441, 444, 445, 455, 456, 458, 477, 481, 484, 488, 491, 498, 523, 526, 527, 531, 532, 533, 206, 306, 168, 164]. Review [270, 339, 46, 173, 278, 376, 249, 310, 381, 454, 494, 236, 349, 244, 374, 100, 86, 150, 179, 108, 114, 19, 450, 377, 175, 53, 177, 29, 136, 232, 305, 410, 446, 497, 82, 345, 346, 370, 420, 105, 312, 495, 350, 383, 178, 216, 217, 87, 238, 117, 17, 489, 59, 213, 375, 22, 118, 442, 496, 522, 23, 111, 134, 58, 364, 18, 499, 61, 52, 139, 202, 212, 479, 77, 330, 334, 388, 424, 485, 490, 31, 182, 492, 457, 272, 338, 451, 101, 538, 279, 386, 132, 197, 247]. Review [251, 453, 478, 480, 102, 137, 163, 368, 419, 449, 486, 28, 104, 303, 316, 181, 331, 413, 440, 333, 409, 425, 542, 439, 501, 33, 49, 221, 309, 176, 203, 525, 32, 75, 166, 401, 528, 112, 210, 21, 138, 180, 500, 269, 131, 50, 314, 201, 300, 184]. Reviews [250]. Revisit [533]. Riani [297]. Richard [399]. Richtsmeier [345]. Rios [408, 373]. Ripley [315, 541, 316]. Risk [56, 141, 415, 438, 228, 343, 216]. Rob [517]. Robert [115, 422]. Roberts [488]. Robinson [246]. Robust [356, 42, 211, 5, 297]. Robustness [500]. Rohwer [439]. Role [224]. Rolski [29]. Rose [102]. Rosenblatt [538]. Ross [149, 247]. Round [66]. Rounded [391]. Rowe [366]. Roy [19]. Royal [390, 229]. Royall [116, 353]. Rozanov [30]. Ruben [362]. Rudd [195]. Rules [141, 287]. Russia [270]. Rutherford [380]. Ryan [178]. Rychert [22].

S [483, 308, 352, 537, 304, 482, 340, 149, 447, 141, 382, 135, 113, 337, 342, 408, 110, 83, 325, 421, 529, 51, 71, 418, 85, 524, 60, 336, 185, 307, 351, 373, 379, 535, 406, 484, 488, 523, 531, 533, 168, 278, 349, 86, 114, 305, 446, 345, 346, 105, 216, 489, 442, 496, 134, 61, 52, 202, 77, 338, 132, 197, 247, 453, 480, 163, 425, 176, 525, 166, 401, 528, 210, 138, 500, 314, 201, 308, 447, 315, 541, 349, 495, 490, 542, 210, 316]. S-PLUS [308, 315, 490, 447, 349, 495, 542, 210]. S.-C [135]. S.-L [418]. Saad [420]. Sabin [209]. Safety [191]. Sahai [277]. Salkind [214]. Salmon [242]. Same [503]. Sample [26, 263, 192, 2, 509, 157, 38, 282, 256, 225, 427, 4, 28, 358, 226, 477]. Samples [472, 357, 505, 181]. Sampling [192, 353, 174]. Sandifer [208]. Sarabia [131]. Sarmukaddam [489]. SAS [221, 396, 321, 383]. Saunders [326]. Scan [143, 531]. Schaefer [48]. Schechter [86]. Scheines [384]. Scheme [39]. Schemes [156]. Schimek [423]. Schlaifer [211]. Schmidli [29]. Schmidt [29]. Schnabel [146]. Schools [504, 193]. Schuerman [241]. Schumacker [495]. Schwarz [86]. Science [42, 378, 273, 458]. Sciences [157, 83, 432, 207, 523, 168, 100, 450, 377, 497, 17, 485, 492, 538]. Scientist [27]. Scientists [379, 53, 335, 383, 23, 424]. Scoring [386]. Search [127, 148, 384]. Searle [446]. Seber [172]. Second [492]. Sectional [192]. Sekaran [208]. Selected [225, 181]. Selection [94, 95, 267, 18]. Self [364, 181]. Self-Instruction [364]. Self-Selected [181]. Semimartingales [85]. Seminar [113]. Seneta [411]. Senn [496]. Sense [364]. Sensing [179]. Separating [319]. Sequences [293]. Sequential [81, 298, 527]. Serfling [454]. Serfozo [215]. Series [352, 537, 235, 276, 530, 278, 542, 221, 84, 375, 499, 425, 299]. Serre [521]. Session [97, 230, 359, 513]. Set [144]. Set-Indexed [144]. Settings [204]. Several [65]. Severini [381]. Shaded [32]. Shape [82, 82]. Shapes [345]. Shaw [239, 418]. Sheskin [455]. Shevlin [243]. Shifts [428].

Shiryaev [346]. Shorack [456]. Should [36, 415]. Shumway [278]. Signal [537, 407]. Siguaw [334]. Sijtsma [457]. Silverman [313]. Simple [38, 474]. Simplest [258]. Simplicity [501]. Simpson [51, 77]. Simulation [341, 65, 484, 257, 534, 241, 218]. Simulation-Based [241]. Simulations [344, 396]. Sincich [172]. Single [387]. Single-Case [387]. Singpurwalla [216]. Sir [12]. Sircar [271]. Sirken [86]. Size [2, 136, 4, 358, 226]. Sizes [509, 427]. Skiena [382]. Skills [342, 139]. Small [26, 387, 192, 275, 357, 282, 303]. Small- [387]. Smith [407, 266, 327, 239, 22]. Smoothing [154, 423]. Snedecor [263]. Snijders [217]. Soccer [290, 69, 68]. Social [157, 27, 24, 335, 168, 100, 432, 232, 497, 383, 23, 479]. Society [390, 51, 77, 102, 229]. Sociology [374]. Sofroniou [27]. Software [152, 216]. Solutions [487, 163]. Solving [415]. Some [504, 310]. Sorting [137]. Sources [449]. Space [102]. Spaces [370]. Spatial [122, 49, 351, 236, 179]. Speaking [106]. Specific [146]. Specification [131]. Spector [383]. Speed [244]. Spencer [19]. Spirit [503]. Spirtes [384]. Spline [493]. Sport [290]. Sports [473, 44]. Spread [96]. Sprott [458]. SPSS [237, 61, 24, 232, 221, 350]. Spurious [395]. Square [498]. Stage [93]. Stages [38]. Standard [185, 67, 285]. Stata [60, 176]. Statements [320]. Stationarity [219]. Stationary [122]. Statistical [1, 57, 16, 153, 344, 88, 25, 83, 115, 55, 390, 85, 540, 60, 211, 307, 445, 455, 491, 164, 494, 345, 495, 229, 499, 388, 538, 176, 131, 26, 116, 385, 408, 76, 273, 366, 524, 140, 367, 373, 458, 532, 114, 377, 178, 216, 87, 58, 139, 492, 331, 440]. Statistically [106]. Statistician [255, 467, 224, 19]. Statisticians [110, 456, 411]. Statistics [237, 447, 113, 460, 27, 315, 335, 504, 228, 209, 143, 185, 541, 289, 310, 349, 244, 63, 179, 193, 497, 17, 111, 52, 35, 424, 485, 101, 50, 464, 462, 536, 141, 135, 170, 147, 234, 103, 214, 301, 535, 406, 444, 531, 249, 381, 454, 175, 177, 117, 213, 31, 449, 109, 72, 172, 378, 372, 51, 207, 214, 403, 498, 232, 346, 522, 77, 528]. Statlabs [244]. Stein [179]. Steps [462]. Stevens [497, 117]. Stigler [500]. Stochastic [448, 271, 30, 215, 165, 480, 47, 29, 132]. Stoffer [278]. Storch [385, 87]. Story [481]. Strata [38]. Strategies [26, 246, 410]. Strategy [288, 270, 8]. Structural [321, 212, 376, 330]. Structures [510]. Stud [288]. Studden [50]. Student [95, 118, 263, 258, 474]. Studies [153, 2, 447, 509]. Study [119, 320, 91, 89, 228]. Subjective [510]. Subjectivity [379]. Substantially [388]. Sugihara [351]. Sum [39, 428]. Surfaces [32]. Survey [157, 20, 477, 86, 464, 486]. Surveys [319, 145]. Survival [248, 75]. Swanson [409]. Swartz [371]. Swinscow [498]. Sydney [283]. Symbolic [41]. Symmetry [39]. System [158]. Systems [148, 510, 418, 311].

T [248, 448, 443, 482, 407, 15, 47, 172, 205, 241, 110, 421, 371, 438, 146, 208, 84, 498, 533, 381, 29, 82, 345, 312, 178, 217, 22, 52, 212, 132, 102]. Table [158]. Table-Tennis [158]. Tables [185, 66, 188, 6]. Tackling [245]. Taking [71, 274]. Taniguchi [499]. Tanur [86, 379]. Target [507]. Teaching [504, 113]. Teams [512, 44]. Techniques [237, 304, 439]. Technological [415]. Temporal [521]. Tennis [158]. Tessellations [351]. Test [293, 152, 262, 386]. Testing [356, 322, 65, 279, 539]. Tests [323, 395, 452, 256, 393, 387, 55]. Tetley [12]. Teugels [29]. Teukolsky [421]. Thatcher [150]. Their [240, 85]. Theorems [454]. Theoretic [18, 124]. Theory [268, 298, 267, 307, 306, 499, 202, 132, 166, 50, 300, 266, 408, 211, 302, 373, 441, 374, 82, 457, 338, 28, 103]. Therneau [248]. Thesaurus [183]. Thiele [535]. Think [214, 175]. Thinking [532]. Thissen [386]. Thode [539]. Thompson [218]. Thorisson [219]. Three [80]. Three-Dimensional [80]. Tiao [352, 500]. Time [352, 537, 2, 235, 276, 530, 84, 375, 499, 102, 425, 542, 221, 32, 299, 341, 278]. Time-Series [537, 299]. Time-Space [102]. Titterington [459, 109]. Todman [387]. Tool [128]. Topics [149, 247]. Torrence [118, 118]. Tourangeau [86]. Tournaments [285]. Townend [424]. Townsend [339]. Transformation [511]. Transformed [102]. Translating [163]. Transportation [418]. Tranter [180]. Trial [36]. Trials [496, 358, 81, 482, 405]. Triangle [481]. Triangular [260]. Tsay [352, 425]. Tukey [294, 211]. Turkington [540]. Turnbull [81]. Tweedie [399]. Twice [468]. Two [256, 393, 66, 427, 80]. Two- [80]. Two-Sample [256]. Tye [113].

U [47, 146, 208]. UK [449, 91, 319, 506]. Umfassendes [461]. Uncertainty [429, 290, 49]. Understanding [211, 495]. Unified [238]. Uniformly [39]. Uninitiated [334]. Union [320]. Univariate [392, 183]. Update [83]. Updating [506]. Use [356, 350, 5, 358]. User [312, 309]. Uses [48]. Using [237, 27, 366, 24, 512, 60, 473, 494, 495, 321, 3, 92, 176, 61]. Utting [295].

V [16, 412, 147, 477, 498, 150, 29, 350].

Vaart [249, 31]. Vague [356]. Valentin [46]. Validation [108]. Validity [292]. Valliant [353]. Value [128, 307]. Values [307]. Vardeman [314]. Variables [2, 506]. Variance [277, 94, 318, 358]. Varying [122]. Vassilicos [313]. Vaupel [32, 150, 108]. vCJD [76]. Venables [315, 541, 316]. Venning [264]. Version [312]. versus [188]. Vetterling [421]. via [320, 371, 251].

Victory [287]. Viel [56]. Vining [450]. Voelkl [61]. vol [461]. Volatility [271]. Vols [346]. Volume [120, 252, 389, 543, 83]. Voronoi [351]. Vukovich [102].

W [109, 81, 340, 407, 313, 385, 198, 48, 337, 315, 80, 421, 341, 294, 348, 211, 301, 481, 249, 150, 108, 305, 312, 350, 87, 479, 31, 457, 316, 409, 309, 32, 500, 50]. Wahrscheinlichkeitsrechnung [461]. Wainer [279, 386, 181]. Wakefield [236]. Walden [407, 276]. Wallenstein [531]. Wang [542]. Water [91, 506]. Watson [409]. Wavelet [1, 276]. Wavelets [313]. Weakest [433]. Weale [304]. Weaver [460]. Weeks [241]. Weibull [37]. Weichselberger [461]. Weisberg [105]. Well [468]. Well-Known [468]. West [523]. Where [245]. Whitaker [532]. Whittle [251]. Who [214, 283, 71]. Wilcox [388]. Wild [172]. Wilfrid [98]. William [360]. Williams [182]. Wilson [216]. Wimmer [183]. Win [382]. Windows [237, 350, 61]. Windows^{T M} [24]. Winkelmann [220]. Winning [274]. Wisdom [393]. Wishart [519]. without [460]. Wolfe [25]. Wolstenholme [88]. Won [283]. Wood [138]. Woodhouse [312]. Woodward [119, 89]. Work [43, 466]. Wright [462]. Wu [184]. Wykes [52]. Wynn [148]. Wynne [518].

XploRe(R) [340].

Yaffee [221]. Yang [312, 310]. Yao [478]. Yashin [32]. Years [504, 459]. Young [407]. Yu [30]. Yuille [80].

Z [340, 364]. Zeger [442, 525]. Zellner [501]. Zelterman [33]. Zero [540]. Zhenglian [32]. Zhigljavsky [148]. Zivot [542]. Zweirs [87]. Zwiers [385]. Zwillinger [185].

References

Abramovich:2000:WAS

[1] Felix Abramovich, Trevor C. Bailey, and Theofanis Sapatinas. Wavelet analysis and its statistical applications. Journal of the Royal Statistical Society. Series D (The Statistician), 49(1):1–29, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                             URL http://www.jstor.org/stable/
                                             2681252.

Bernardo:2000:SSC

[2] M. V. Patricia Bernardo, Stuart R. Lipsitz, David P. Harrington, and Paul J. Catalano. Sample size calculations for failure time random variables in nonrandomized studies. Journal of the Royal Statistical Society. Series D (The Statistician), 49(1):31–40, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                             www.jstor.org/stable/2681253.

Rigby:2000:CRC

[3] R. A. Rigby and D. M. Stasinopoulos. Construction of reference centiles using mean and dispersion additive models. Journal of the Royal Statistical Society. Series D (The Statistician), 49 (1):41–50, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                             tronic). URL http://www.jstor.org/
                                             stable/2681254.

Soliman:2000:BPP

[4] Ahmed A. Soliman. Bayes prediction in a Pareto lifetime model with random sample size. Journal of the Royal Statistical Society. Series D (The Statistician), 49(1):51–62, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2681255.

Thomas:2000:UBR

[5] G. E. Thomas. Use of the bootstrap in robust estimation of location. Journal of the Royal Statistical Society. Series D (The Statistician), 49(1): 63–77, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681256.

Upton:2000:CDM

[6] Graham J. G. Upton. Cobweb diagrams for multiway contingency tables. Journal of the Royal Statistical Society. Series D (The Statistician), 49 (1):79–85, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681257.

deLuna:2000:PIB

[7] Xavier de Luna. Prediction intervals based on autoregression forecasts. Journal of the Royal Statistical Society. Series D (The Statistician), 49(1): 87–93, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681258.

Preston:2000:BSL

[8] Ian Preston and Jonathan Thomas. Batting strategy in limited overs Cricket. Journal of the Royal Statistical Society. Series D (The Statistician), 49(1): 95–106, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681259.

Clarke:2000:C

[9] Geoff Clarke, J. A. Nelder, and Douglas C. Montgomery. Correspondence. Journal of the Royal Statistical Society. Series D (The Statistician), 49(1): 107–110, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2681260.

Mallows:2000:CD

[10] Colin Mallows. Cuthbert Daniel, 1905– 1997. Journal of the Royal Statistical Society. Series D (The Statistician), 49(1):111, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2681261.

Liddell:2000:BPE

[11] Douglas Liddell. Brian Patrick emmett. Journal of the Royal Statistical Society. Series D (The Statistician), 49(1): 111–113, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2681262.

Moore:2000:SHT

[12] Peter G. Moore. Sir Herbert Tetley, 1908–1999. Journal of the Royal Statistical Society. Series D (The Statistician), 49(1):113–114, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2681263.

Anonymous:2000:FMa

[13] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 49(1):115, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2681251.

Collins:2000:BRB

[14] Sean Collins. Book review: Constrained Markov Decision Processes, by E. Altman. Journal of the Royal Statistical Society. Series D (The Statistician), 49(1):116, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681264.

Brooks:2000:BRB

[15] Stephen Brooks. Book review: Encyclopedia of Biostatistics, by P. Armitage; T. Colton. Journal of the Royal Statistical Society. Series D (The Statistician), 49(1):116–117, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681265.

Auton:2000:BRBa

[16] Tim Auton. Book review: Comparative Statistical Inference, by V. Barnett. Journal of the Royal Statistical Society. Series D (The Statistician), 49 (1):117–118, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681266.

Penenberg:2000:BRB

[17] Darryl Neil Penenberg. Book review: Applied Multivariate Statistics in Geohydrology and Related Sciences, by C. E. Brown. Journal of the Royal Statistical Society. Series D (The Statistician), 49 (1):118–119, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681267.

Sapatinas:2000:BRB

[18] Theofanis Sapatinas. Book review: Model Selection and Inference: A Practical Information-Theoretic Approach, by K. P. Burnham; D. R. Anderson. Journal of the Royal Statistical Society. Series D (The Statistician), 49(1): 119–120, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681268.

Mayo:2000:BRB

[19] Oliver Mayo. Book review: Roy Geary 1896–1983, Irish Statistician: Centenary Lecture by John E. Spencer and Associated Papers, by D. Conniffe. Journal of the Royal Statistical Society. Series D (The Statistician), 49(1):120–121, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2681269.

Collop:2000:BRB

[20] Michael Collop. Book review: Computer Assisted Survey Information Collection, by M. P. Couper. Journal of the Royal Statistical Society. Series D (The Statistician), 49(1):121–122, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2681270.

Woodward:2000:BRB

[21] P. W. Woodward. Book review: Fractional Factorial Plans, by A. Dey; R. Mukerjee. Journal of the Royal Statistical Society. Series D (The Statistician), 49(1):122, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681271.

Reed:2000:BRBa

[22] Matt Reed. Book review: Maximum Entropy and Bayesian Methods, by G. J. Erickson; J. T. Rychert; C. R. Smith. Journal of the Royal Statistical Society. Series D (The Statistician), 49(1): 122–123, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681272.

Robertson:2000:BRB

[23] Chris Robertson. Book review: Neural Networks: An Introductory Guide for Social Scientists, by G. D. Garson. Journal of the Royal Statistical Society. Series D (The Statistician), 49(1): 123–124, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681273.

Hemmett:2000:BRBa

[24] Lindsay Hemmett. Book review: Exploring Social Issues Using SPSS(R) for Windows^{T M}, by J. F. Healey; E. Babbie; F. Halley. Journal of the Royal Statistical Society. Series D (The Statistician), 49(1):124–125, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681274.

Coolen:2000:BRB

[25] Frank Coolen. Book review: Nonparametric Statistical Methods, by M. Hollander; D. A. Wolfe. Journal of the Royal Statistical Society. Series D (The Statistician), 49(1):125–126, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2681275.

Auton:2000:BRBb

[26] Tim Auton. Book review: Statistical Strategies for Small Sample Research, by R. H. Hoyle. Journal of the Royal Statistical Society. Series D (The Statistician), 49(1):126–127, ???? 2000. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2681276.

Espejo:2000:BRBa

[27] Mariano Ruiz Espejo. Book review: The Multivariate Social Scientist: Introductory Statistics Using Generalized Linear Models, by G. D. Hutcheson; N. Sofroniou. Journal of the Royal Statistical Society. Series D (The Statistician), 49(1):127, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884

                                        (electronic). URL http://www.jstor.
                                        org/stable/2681277.

Thas:2000:BRB

[28] Olivier Thas. Book review: Elements of Large-Sample Theory, by E. L. Lehmann. Journal of the Royal Statistical Society. Series D (The Statistician), 49(1):127–128, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2681278.

Minkova:2000:BRB

[29] Leda Minkova. Book review: Stochastic Processes for Insurance and Finance, by T. Rolski; H. Schmidli; V. Schmidt; J. Teugels. Journal of the Royal Statistical Society. Series D (The Statistician), 49(1):128–129, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2681279.

Gani:2000:BRBa

[30] Joe Gani. Book review: Random Fields and Stochastic Partial Differential Equations, by Yu. A. Rozanov. Journal of the Royal Statistical Society. Series D (The Statistician), 49 (1):130, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681280.

Skouras:2000:BRB

[31] Kostas Skouras. Book review: Asymptotic Statistics, by A. W. van der Vaart. Journal of the Royal Statistical Society. Series D (The Statistician), 49(1): 130–131, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681281.

Wang:2000:BRBa

[32] Duolao Wang. Book review: Population Data at a Glance: Shaded Contour Maps of Demographic Surfaces over Age and Time, by J. W. Vaupel; W. Zhenglian; K. F. Andreev; A. I. Yashin. Journal of the Royal Statistical Society. Series D (The Statistician), 49(1): 131–132, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681282.

Ugarte:2000:BRBa

[33] M. Dolores Ugarte. Book review: Models for Discrete Data, by D. Zelterman. Journal of the Royal Statistical Society. Series D (The Statistician), 49(1): 132–133, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681283.

Anonymous:2000:BMa

[34] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 49(1):??, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2681284.

Senn:2000:CCP

[35] Stephen Senn. Consensus and controversy in pharmaceutical statistics. Journal of the Royal Statistical Society. Series D (The Statistician), 49(2): 135–176, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2680966.

Gittins:2000:HLS

[36] John Gittins and Hamid Pezeshk. How large should a clinical trial be? Journal of the Royal Statistical Society. Series D (The Statistician), 49(2):177–187, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2680967.

Davison:2000:IPW

[37] A. C. Davison and F. Louzada-Neto. Inference for the poly-Weibull model. Journal of the Royal Statistical Society. Series D (The Statistician), 49(2): 189–196, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2680968.

Clark:2000:OAS

[38] Robert G. Clark and D. G. Steel. Optimum allocation of sample to strata and stages with simple additional constraints. Journal of the Royal Statistical Society. Series D (The Statistician), 49 (2):197–207, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2680969.

Atienza:2000:UMP

[39] O. O. Atienza, L. C. Tang, and B. W. Ang. A uniformly most powerful cumulative sum scheme based on symmetry. Journal of the Royal Statistical Society. Series D (The Statistician), 49 (2):209–217, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2680970.

Chakraborti:2000:PPP

[40] S. Chakraborti and P. van der Laan. Precedence probability and prediction intervals. Journal of the Royal Statistical Society. Series D (The Statistician), 49(2):219–228, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2680971.

Rose:2000:SML

[41] Colin Rose and Murray D. Smith. Symbolic maximum likelihood estimation with Mathematica. Journal of the Royal Statistical Society. Series D (The Statistician), 49(2):229–240, ???? 2000. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2680972.

Deniz:2000:RBP

[42] E. Gómez Déniz, F. J. Vázquez Polo, and A. Hernández Bastida. Robust Bayesian premium principles in actuarial science. Journal of the Royal Statistical Society. Series D (The Statistician), 49(2):241–252, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2680973.

Forrest:2000:MRW

[43] David Forrest and Robert Simmons. Making up the results: The work of the Football Pools Panel, 1963–1997. Journal of the Royal Statistical Society. Series D (The Statistician), 49(2): 253–260, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2680974.

Knorr-Held:2000:DRS

[44] Leonhard Knorr-Held. Dynamic rating of sports teams. Journal of the Royal Statistical Society. Series D (The Statistician), 49(2):261–276, ???? 2000. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2680975.

Anonymous:2000:FMb

[45] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 49(2):277, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2680965.

Kumar:2000:BRB

[46] Kuldeep Kumar. Book review: Neural Networks, by H. Abdi; D. Valentin; B. Edelman. Journal of the Royal Statistical Society. Series D (The Statistician), 49(2):278, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884

                                        (electronic). URL http://www.jstor.
                                        org/stable/2680976.

Burrell:2000:BRBa

[47] Quentin Burrell. Book review: Stochastic Models in Reliability, by T. Aven; U. Jensen. Journal of the Royal Statistical Society. Series D (The Statistician), 49(2):278–279, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2680977.

Dialsingh:2000:BRBa

[48] Isaac Dialsingh. Book review: The Uses and Misuses of Data and Models, by W. J. Bradley; C. K. Schaefer. Journal of the Royal Statistical Society. Series D (The Statistician), 49 (2):279–280, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2680978.

Ugarte:2000:BRBb

[49] M. Dolores Ugarte. Book review: Geostatistics-Modeling Spatial Uncertainty, by J. P. Chiles; P. Delfiner. Journal of the Royal Statistical Society. Series D (The Statistician), 49(2): 280–281, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2680979.

Young:2000:BRB

[50] Philip Young. Book review: The Theory of Canonical Moments with Applications in Statistics, Probability, and Analysis, by H. Dette; W. J. Studden. Journal of the Royal Statistical Society. Series D (The Statistician), 49(2): 281–282, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2680980.

Hewson:2000:BRBa

[51] Paul Hewson. Book review: Statistics in Society: The Arithmetic of Politics, by D. Dorling; S. Simpson. Journal of the Royal Statistical Society. Series D (The Statistician), 49(2):282–283, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2680981.

Scott:2000:BRB

[52] Marian Scott. Book review: A Dictionary of Statistics for Psychologists, by B. S. Everitt; T. Wykes. Journal of the Royal Statistical Society. Series D (The Statistician), 49(2):283–284, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2680982.

Michailidis:2000:BRB

[53] George Michailidis. Book review: Essential MATLAB for Scientists and Engineers, by B. D. Hahn. Journal of the Royal Statistical Society. Series D (The Statistician), 49(2):284, ???? 2000. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2680983.

Haslett:2000:BRB

[54] John Haslett. Book review: Learning in Graphical Models, by M. I. Jordon. Journal of the Royal Statistical Society. Series D (The Statistician), 49 (2):285, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2680984.

Gray:2000:BRB

[55] Alison Gray. Book review: 100 Statistical Tests, by G. K. Kanji. Journal of the Royal Statistical Society. Series D (The Statistician), 49(2):286, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2680985.

Brix:2000:BRB

[56] A. Brix. Book review: Disease Mapping and Risk Assessment for Public Health, by A. Lawson; A. Biggeri; D. Böhning; E. Lesaffre; J.-F. Viel; R. Bertollini. Journal of the Royal Statistical Society. Series D (The Statistician), 49(2): 286–287, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2680986.

Apostolides:2000:BRB

[57] Michael G. Apostolides. Book review: Revealing Statistical Principles, by J. K. Lindsey. Journal of the Royal Statistical Society. Series D (The Statistician), 49(2):287, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2680987.

Roddam:2000:BRB

[58] Andrew W. Roddam. Book review: Statistical Analysis of Categorical Data, by C. J. Lloyd. Journal of the Royal Statistical Society. Series D (The Statistician), 49(2):287–288, ???? 2000. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2680988.

Pewsey:2000:BRB

[59] Arthur Pewsey. Book review: Generalized Method of Moments Estimation, by L. Mátyás. Journal of the Royal Statistical Society. Series D (The Statistician), 49(2):288–289, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2680989.

Jeffs:2000:BRB

[60] James Jeffs. Book review: A Handbook of Statistical Analyses Using Stata, by S. Rabe-Hesketh; B. Everitt. Journal of the Royal Statistical Society. Series D (The Statistician), 49(2):290, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2680990.

Sarker:2000:BRB

[61] Shah Jalal Sarker. Book review: Using SPSS for Windows — Data Analysis and Graphics, by K. E. Voelkl; S. B. Gerber. Journal of the Royal Statistical Society. Series D (The Statistician), 49(2):290–291, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2680991.

Anonymous:2000:BMb

[62] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 49(2):??, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2680992.

Lindley:2000:PS

[63] Dennis V. Lindley. The philosophy of statistics. Journal of the Royal Statistical Society. Series D (The Statistician), 49(3):293–337, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2681060.

Fan:2000:BMP

[64] Y. Fan and S. P. Brooks. Bayesian modelling of prehistoric corbelled domes. Journal of the Royal Statistical Society. Series D (The Statistician), 49(3): 339–354, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681061.

Karlis:2000:SCS

[65] Dimitris Karlis and Evdokia Xekalaki. A simulation comparison of several procedures for testing the Poisson assumption. Journal of the Royal Statistical Society. Series D (The Statistician), 49 (3):355–382, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681062.

Nelder:2000:ACT

[66] John A. Nelder. The analysis of contingency tables with one factor as the response: Round two. Journal of the Royal Statistical Society. Series D (The Statistician), 49(3):383–388, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2681063.

Longford:2000:ESE

[67] Nicholas T. Longford. On estimating standard errors in multilevel analysis. Journal of the Royal Statistical Society. Series D (The Statistician), 49(3): 389–398, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681064.

Rue:2000:PRA

[68] Håvard Rue and Øyvind Salvesen. Prediction and retrospective analysis of soccer matches in a league. Journal of the Royal Statistical Society. Series D (The Statistician), 49(3):399–418, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2681065.

Koning:2000:BCD

[69] Ruud H. Koning. Balance in competition in Dutch soccer. Journal of the Royal Statistical Society. Series D (The Statistician), 49(3):419–431, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2681066.

Anonymous:2000:FMc

[70] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 49(3):433, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2681059.

Hewson:2000:BRBb

[71] Paul Hewson. Book review: Who Counts?: The Politics of Census-Taking in Contemporary America, by M. J. Anderson; S. E. Fienberg. Journal of the Royal Statistical Society. Series D (The Statistician), 49(3):434, ???? 2000. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2681067.

Arteche:2000:BRB

[72] Josu Arteche. Book review: Statistics for Long-Memory Processes, by J. Beran. Journal of the Royal Statistical Society. Series D (The Statistician), 49 (3):434–436, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681068.

Crowder:2000:BRB

[73] Martin Crowder. Book review: Convergence of Probability Measures, by P. Billingsley. Journal of the Royal Statistical Society. Series D (The Statistician), 49(3):436–437, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681069.

Harris:2000:BRB

[74] Ray Harris. Book review: Introduction to Multivariate Analysis, by C. Chatfield; A. J. Collins. Journal of the Royal Statistical Society. Series D (The Statistician), 49(3):437, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2681070.

Wang:2000:BRBb

[75] Duolao Wang. Book review: Modelling Survival Data in Medical Research, by D. Collett. Journal of the Royal Statistical Society. Series D (The Statistician), 49(3):437–438, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681071.

Greenhalgh:2000:BRB

[76] David Greenhalgh. Book review: Statistical Aspects of BSE and vCJD: Models for Epidemics, by C. A. Donnelly; N. M. Ferguson. Journal of the Royal Statistical Society. Series D (The Statistician), 49(3):438–439, ???? 2000. CODEN ????

ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2681072.

Shade:2000:BRB

[77] John Shade. Book review: Statistics in Society: The Arithmetic of Politics, by D. Dorling; S. Simpson. Journal of the Royal Statistical Society. Series D (The Statistician), 49(3):439–440, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2681073.

Kemp:2000:BRBa

[78] A. W. Kemp. Book review: Resampling Methods, by P. I. Good. Journal of the Royal Statistical Society. Series D (The Statistician), 49(3):440–441, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2681074.

Auton:2000:BRBc

[79] Tim Auton. Book review: Classification, by A. D. Gordon. Journal of the Royal Statistical Society. Series D (The Statistician), 49(3):441–442, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2681075.

Gani:2000:BRBb

[80] Joe Gani. Book review: Two- and Three-Dimensional Patterns of the Face, by P. W. Hallinan; G. G. Gordon; A. L. Yuille; P. Giblin; D. Mumford. Journal of the Royal Statistical Society. Series D (The Statistician), 49(3): 442–444, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2681076.

Auton:2000:BRBd

[81] Tim Auton. Book review: Group Sequential Methods with Applications to Clinical Trials, by C. Jennison; B. W. Turnbull. Journal of the Royal Statistical Society. Series D (The Statistician), 49(3):444–445, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681077.

Molchanov:2000:BRB

[82] Ilya Molchanov. Book review: Shape and Shape Theory, by D. G. Kendall; D. Barden; T. K. Carne; H. Le. Journal of the Royal Statistical Society. Series D (The Statistician), 49(3):445–446, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2681078.

Gani:2000:BRBc

[83] Joe Gani. Book review: Encyclopedia of Statistical Sciences Update Volume 3, by S. Kotz; C. B. Read; D. L. Banks. Journal of the Royal Statistical Society. Series D (The Statistician), 49 (3):446–448, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681079.

Kemp:2000:BRBb

[84] Malcolm Kemp. Book review: The Econometric Modelling of Financial Time Series, by T. C. Mills. Journal of the Royal Statistical Society. Series D (The Statistician), 49(3):448, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2681080.

Heyde:2000:BRB

[85] C. C. Heyde. Book review: Semimartingales and Their Statistical Inference, by B. L. S. Prakasa Rao. Journal of the Royal Statistical Society. Series D (The Statistician), 49(3):449, ???? 2000. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2681081.

Lynn:2000:BRB

[86] Peter Lynn. Book review: Cognition and Survey Research, by M. G. Sirken; D. J. Herrmann; S. Schechter; N. Schwarz; J. M. Tanur; R. Tourangeau. Journal of the Royal Statistical Society. Series D (The Statistician), 49(3):449–450, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2681082.

Neykov:2000:BRB

[87] Neyko M. Neykov. Book review: Statistical Analysis in Climate Research, by H. von Storch; F. W. Zweirs. Journal of the Royal Statistical Society. Series D (The Statistician), 49(3):450–451, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2681083.

Burrell:2000:BRBb

[88] Quentin Burrell. Book review: Reliability Modelling: A Statistical Approach, by L. C. Wolstenholme. Journal of the Royal Statistical Society. Series D (The Statistician), 49(3):451–452, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2681084.

Hemmett:2000:BRBb

[89] Lindsay Hemmett. Book review: Epidemiology: Study Design and Data Analysis, by M. Woodward. Journal of the Royal Statistical Society. Series D (The Statistician), 49(3):452–453, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2681085.

Anonymous:2000:BMc

[90] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 49(3):??, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2681086.

Garthwaite:2000:QEO

[91] Paul H. Garthwaite and Anthony O’Hagan. Quantifying expert opinion in the UK water industry: An experimental study. Journal of the Royal Statistical Society. Series D (The Statistician), 49(4):455–477, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681030.

Stasinopoulos:2000:MRG

[92] D. M. Stasinopoulos, R. A. Rigby, and L. Fahrmeir. Modelling rental guide data using mean and dispersion additive models. Journal of the Royal Statistical Society. Series D (The Statistician), 49(4):479–493, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681031.

Chan:2000:ISP

[93] Jennifer S. K. Chan. Initial stage problem in autoregressive binary regression. Journal of the Royal Statistical Society. Series D (The Statistician), 49(4): 495–502, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2681032.

Bertolino:2000:BMS

[94] Francesco Bertolino, Walter Racugno, and Elías Moreno. Bayesian model selection approach to analysis of variance under heteroscedasticity. Journal of the Royal Statistical Society. Series D (The Statistician), 49(4):503–517, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2681033.

Boland:2000:SSR

[95] Philip J. Boland and Kevin Hutchinson. Student selection of random digits. Journal of the Royal Statistical Society. Series D (The Statistician), 49 (4):519–529, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2681034.

Haigh:2000:KCB

[96] John Haigh. The Kelly criterion and bet comparisons in spread betting. Journal of the Royal Statistical Society. Series D (The Statistician), 49(4):531–539, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2681035.

Anonymous:2000:RCS

[97] Anonymous. Report of the Council for the session 1999–2000. Journal of the Royal Statistical Society. Series D (The Statistician), 49(4):541–613, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2681036.

Bissell:2000:WAP

[98] Derek Bissell. Wilfrid Alan Pridmore, 1921–2000. Journal of the Royal Statistical Society. Series D (The Statistician), 49(4):615–617, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681037.

Anonymous:2000:FMd

[99] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 49(4):619, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2681029.

Longford:2000:BRB

[100] Nick Longford. Book review: Research Methodology in the Social, Behavioural and Life Sciences, by H. J. Adèr; G. J. Mellenbergh. Journal of the Royal Statistical Society. Series D (The Statistician), 49(4):620–621, ???? 2000. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2681038.

Stallard:2000:BRB

[101] Nigel Stallard. Book review: Practical Statistics for Medical Research, by D. G. Altman. Journal of the Royal Statistical Society. Series D (The Statistician), 49(4):621, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2681039.

Swires-Hennessy:2000:BRB

[102] Ed Swires-Hennessy. Book review: A Society Transformed: Hungary in Time-Space Perspective, by R. Andorka; T. Kolosi; R. Rose; G. Vukovich. Journal of the Royal Statistical Society. Series D (The Statistician), 49(4):621–622, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2681040.

Jones:2000:BRB

[103] M. C. Jones. Book review: Theory of Multivariate Statistics, by M. Bilodeau; D. Brenner. Journal of the Royal Statistical Society. Series D (The Statistician), 49(4):622–623, ???? 2000. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2681041.

Thomas:2000:BRB

[104] Len Thomas. Book review: Bootstrap Methods: A Practitioner’s Guide, by M. R. Chernick. Journal of the Royal Statistical Society. Series D (The Statistician), 49(4):623–624, ???? 2000. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2681042.

Moreno:2000:BRB

[105] Victor Moreno. Book review: Applied Regression Including Computing and Graphics, by R. D. Cook; S. Weisberg. Journal of the Royal Statistical Society. Series D (The Statistician), 49 (4):624–625, ???? 2000. CODEN ????

ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681043.

Kemp:2000:BRBc

[106] A. W. Kemp. Book review: “Statistically Speaking” — A Dictionary of Quotations, by C. C. Gaither; A. E. Cavazos-Gaither. Journal of the Royal Statistical Society. Series D (The Statistician), 49 (4):625–626, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681044.

Espejo:2000:BRBb

[107] Mariano Ruiz Espejo. Book review:

Analyzing Quantitative Accounts, by R. Hodson. Journal of the Royal Statistical Society. Series D (The Statistician), 49(4):626, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681045.

Matthews:2000:BRBa

[108] Fiona Matthews. Book review: Validation of Exceptional Longevity, by B. Jeune; J. W. Vaupel. Journal of the Royal Statistical Society. Series D (The Statistician), 49(4):627, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2681046.

Aitkin:2000:BRB

[109] Murray Aitkin. Book review: Statistics and Neural Networks: Advances at the Interface, by J. W. Kay; D. M. Titterington. Journal of the Royal Statistical Society. Series D (The Statistician), 49 (4):627–628, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2681047.

Fernandez:2000:BRB

[110] Carmen Fernández. Book review: Bayesian Methods (An Analysis for Statisticians and Interdisciplinary Researchers), by T. Leonard; J. S. J. Hsu. Journal of the Royal Statistical Society. Series D (The Statistician), 49(4): 628–629, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2681048.

Robinson:2000:BRB

[111] Andrew Robinson. Book review: Relating Statistics and Experimental Design: An Introduction, by I. P. Levin. Journal of the Royal Statistical Society. Series D (The Statistician), 49(4): 629–630, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2681049.

Wood:2000:BRB

[112] Simon Wood. Book review: Models for Repeated Measurements, by J. K. Lindsey. Journal of the Royal Statistical Society. Series D (The Statistician), 49 (4):630–631, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2681050.

Dialsingh:2000:BRBb

[113] Isaac Dialsingh. Book review: Teaching Health Statistics: Lesson and Seminar Outlines, by S. K. Lwanga; C.-Y. Tye; O. Ayeni. Journal of the Royal Statistical Society. Series D (The Statistician), 49(4):631, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2681051.

Matthews:2000:BRBb

[114] Fiona Matthews. Book review: Statistical Inference: An Integrated Approach, by H. S. Migon; D. Gamerman. Journal of the Royal Statistical Society. Series D (The Statistician), 49 (4):631–632, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681052.

Gentle:2000:BRB

[115] James E. Gentle. Book review: Monte Carlo Statistical Methods, by C. P. Robert; G. Casella. Journal of the Royal Statistical Society. Series D (The Statistician), 49(4):632–633, ???? 2000. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2681053.

Chadwick:2000:BRB

[116] Thomas Chadwick. Book review: Statistical Evidence: A Likelihood Paradigm, by R. Royall. Journal of the Royal Statistical Society. Series D (The Statistician), 49(4):633–634, ???? 2000. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2681054.

Oliver:2000:BRB

[117] F. R. Oliver. Book review: Intermediate Statistics, by J. Stevens. Journal of the Royal Statistical Society. Series D (The Statistician), 49(4):634–635, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2681055.

Reed:2000:BRBb

[118] R. J. Reed. Book review: The Student’s Introduction to MATHEMAT-ICA: A Handbook for Precalculus, Calculus and Linear Algebra, by B. F. Torrence; E. A. Torrence. Journal of the Royal Statistical Society. Series D (The Statistician), 49(4):635, ???? 2000. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2681056.

Armitage:2000:BRB

[119] Peter Armitage. Book review: Epidemiology: Study Design and Data Analysis, by M. Woodward. Journal of the Royal Statistical Society. Series D (The Statistician), 49(4):635–636, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2681057.

Anonymous:2000:VI

[120] Anonymous. Volume information. Journal of the Royal Statistical Society. Series D (The Statistician), 49(4):637–641, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2681028.

Anonymous:2000:BMd

[121] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 49(4):??, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2681058.

Dreesman:2001:NSC

[122] Johannes M. Dreesman and Gerhard Tutz. Non-stationary conditional models for spatial data based on varying coefficients. Journal of the Royal Statistical Society. Series D (The Statistician), 50(1):1–15, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681153.

Moerbeek:2001:OED

[123] Mirjam Moerbeek, Gerard J. P. Van Breukelen, and Martijn P. F. Berger. Optimal experimental designs for multilevel logistic models. Journal of the Royal Statistical Society. Series D (The Statistician), 50(1):17–30, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2681154.

Walker:2001:DTA

[124] S. G. Walker, E. Gutiérrez-Peña, and P. Muliere. A decision theoretic approach to model averaging. Journal of the Royal Statistical Society. Series D (The Statistician), 50(1):31–39, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2681155.

Kelly:2001:MLD

[125] Gabrielle E. Kelly. The median lethal dose — design and estimation. Journal of the Royal Statistical Society. Series D (The Statistician), 50(1):41–50, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/

2681156. See corrigendum [222].

Julious:2001:IEC

[126] Steven A. Julious. Inference and estimation in a changepoint regression problem. Journal of the Royal Statistical Society. Series D (The Statistician), 50 (1):51–61, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2681157.

Atkinson:2001:RDB

[127] Anthony C. Atkinson and Marco Riani. Regression diagnostics for binomial data from the forward search. Journal of the Royal Statistical Society. Series D (The Statistician), 50(1):63–78, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2681158.

Berger:2001:VII

[128] Vance W. Berger. The P -value interval as an inferential tool. Journal of the Royal Statistical Society. Series D (The Statistician), 50(1):79–85, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2681159.

Lipsitz:2001:PCC

[129] Stuart R. Lipsitz, Traci Leong, Joseph Ibrahim, and Steven Lipshultz. A partial correlation coefficient and coefficient of determination for multivariate normal repeated measures data. Journal of the Royal Statistical Society. Series D (The Statistician), 50(1):87–95, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2681160.

Anonymous:2001:FMa

[130] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 50(1):97, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2681152.

Xekalaki:2001:BRB

[131] Evdokia Xekalaki. Book review: Conditional Specification of Statistical Models, by B. C. Arnold; E. Castillo; J. M. Sarabia. Journal of the Royal Statistical Society. Series D (The Statistician), 50(1):98–99, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681161.

Stoyanov:2001:BRBa

[132] Jordan Stoyanov. Book review: Stochastic and Differential Games: Theory and Numerical Methods, by M. Bardi; T. E. S. Raghavan; T. Parthasarathy. Journal of the Royal Statistical Society. Series D (The Statistician), 50(1):99, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2681162.

Kemp:2001:BRBa

[133] David Kemp. Book review: What is Random?: Chance and Order in Mathematics and Life, by E. Beltrami. Journal of the Royal Statistical Society. Series D (The Statistician), 50(1): 99–100, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681163.

Robinson:2001:BRB

[134] Andrew Robinson. Book review: Regression Analysis, by Example, by S. Chatterjee; A. S. Hadi; B. Price. Journal of the Royal Statistical Society. Series D (The Statistician), 50(1):101, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2681164.

Day:2001:BRB

[135] Simon Day. Book review: Encyclopedia of Biopharmaceutical Statistics, by S.-C. Chow. Journal of the Royal Statistical Society. Series D (The Statistician), 50 (1):101–103, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2681165.

Mitchell:2001:BRBa

[136] Lada Mitchell. Book review: Effect Size for ANOVA Designs, by J. M. Cortina; H. Nouri. Journal of the Royal Statistical Society. Series D (The Statistician), 50(1):103, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2681166.

Swires-Hennessy:2001:BRBa

[137] Ed Swires-Hennessy. Book review: Sorting Data: Collection and Analysis, by A. P. M. Coxon. Journal of the Royal Statistical Society. Series D (The Statistician), 50(1):103–104, ???? 2001. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2681167.

Woodward:2001:BRBa

[138] P. W. Woodward. Book review: Fitting Equations to Data: Computer Analysis of Multifactor Data, by C. Daniel; F. S. Wood. Journal of the Royal Statistical Society. Series D (The Statistician), 50 (1):104–105, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2681168.

Scott:2001:BRBa

[139] Marian Scott. Book review: Statistical and Data Handling Skills in Biology, by R. Ennos. Journal of the Royal Statistical Society. Series D (The Statistician), 50(1):105–106, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681169.

Kemp:2001:BRBb

[140] A. W. Kemp. Book review: Statistical Distributions, by M. Evans; N. Hastings; B. Peacock. Journal of the Royal Statistical Society. Series D (The Statistician), 50(1):106–107, ???? 2001. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2681170.

Burrell:2001:BRBa

[141] Quentin Burrell. Book review: Chance Rules: An Informal Guide to Probability, Risk and Statistics, by B. S. Everitt. Journal of the Royal Statistical Society. Series D (The Statistician), 50(1): 107–108, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681171.

Kemp:2001:BRBc

[142] Gordon Kemp. Book review: Mathematical Methods and Models for Economists, by A. de la Fuente. Journal of the Royal Statistical Society. Series D (The Statistician), 50(1):108, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2681172.

Kemp:2001:BRBd

[143] A. W. Kemp. Book review: Scan Statistics and Applications, by J. Glaz; N. Balakhrishnan. Journal of the Royal Statistical Society. Series D (The Statistician), 50(1):109, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2681173.

Harris:2001:BRB

[144] Simon Harris. Book review: Set-Indexed Martingales, by G. Ivanoff; E. Merzbach. Journal of the Royal Statistical Society. Series D (The Statistician), 50(1): 109–110, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681174.

Hewson:2001:BRBa

[145] Paul Hewson. Book review: Analysis of Health Surveys, by E. L. Korn; B. I. Graubard. Journal of the Royal Statistical Society. Series D (The Statistician), 50(1):110–111, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681175.

Hewson:2001:BRBb

[146] Paul Hewson. Book review: Modeling Longitudinal and Multilevel Data: Practical Issues, Applied Approaches and Specific Examples, by T. D. Little; K. U. Schnabel; J. Baumert. Journal of the Royal Statistical Society. Series D (The Statistician), 50(1):111–112, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2681176.

Gadsden:2001:BRB

[147] R. J. Gadsden. Book review: Directional Statistics, by K. V. Mardia; P. E. Jupp. Journal of the Royal Statistical Society. Series D (The Statistician), 50 (1):112–113, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681177.

Clark:2001:BRB

[148] Stephen Clark. Book review: Dynamical Search: Applications of Dynamical Systems in Search and Optimization, by L. Pronzato; H. P. Wynn; A. A. Zhigljavsky. Journal of the Royal Statistical Society. Series D (The Statistician), 50 (1):113–114, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681178.

Boshnakov:2001:BRBa

[149] Georgi Boshnakov. Book review: An Introduction to Mathematical Finance: Options and Other Topics, by S. M. Ross. Journal of the Royal Statistical Society. Series D (The Statistician), 50(1):114, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681179.

Macdonald:2001:BRB

[150] Angus S. Macdonald. Book review: The Force of Mortality at Ages 80 to 120, by A. R. Thatcher; V. Kannisto; J. W. Vaupel. Journal of the Royal Statistical Society. Series D (The Statistician), 50(1):115, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681180.

Anonymous:2001:BMa

[151] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 50(1):??, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2681181.

McDaid:2001:DHL

[152] Kevin McDaid and Simon P. Wilson. Deciding how long to test software. Journal of the Royal Statistical Society. Series D (The Statistician), 50(2): 117–134, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2681088.

Bartfay:2001:SII

[153] Emma Bartfay and Allan Donner. Statistical inferences for interobserver agreement studies with nominal outcome data. Journal of the Royal Statistical Society. Series D (The Statistician), 50(2):135–146, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2681089.

Chatfield:2001:NLM

[154] Chris Chatfield, Anne B. Koehler, J. Keith Ord, and Ralph D. Snyder. A new look at models for exponential smoothing. Journal of the Royal Statistical Society. Series D (The Statistician), 50(2):147–159, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2681090.

Conniffe:2001:WMR

[155] Denis Conniffe and John E. Spencer. When moments of ratios are ratios of moments. Journal of the Royal Statistical Society. Series D (The Statistician), 50(2):161–168, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681091.

Luceno:2001:EAE

[156] Alberto Luceño. Effects of adjustment errors on discrete feed-back dead band control schemes. Journal of the Royal Statistical Society. Series D (The Statistician), 50(2):169–177, ???? 2001. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2681092.

Brunt:2001:ASS

[157] L. Brunt. The advent of the sample survey in the social sciences. Journal of the Royal Statistical Society. Series D (The Statistician), 50(2):179–189, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2681093.

Marcus:2001:NTT

[158] David J. Marcus. New table-tennis rating system. Journal of the Royal Statistical Society. Series D (The Statistician), 50(2):191–208, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681094.

Nelder:2001:C

[159] J. A. Nelder. Correspondence. Journal of the Royal Statistical Society. Series D (The Statistician), 50(2):209–211, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2681095.

Whittle:2001:HD

[160] Peter Whittle. Henry daniels. Journal of the Royal Statistical Society. Series D (The Statistician), 50(2):213–214, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2681096.

Kent:2001:EJR

[161] John Kent. Edwin J. Redfern, 1948– 2000. Journal of the Royal Statistical Society. Series D (The Statistician), 50 (2):214–215, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2681097.

Anonymous:2001:FMb

[162] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 50(2):217, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2681087.

Swires-Hennessy:2001:BRBb

[163] Ed Swires-Hennessy. Book review: Translating Questionnaires and Other Research Instruments: Problems and Solutions, by O. Behling; K. S. Law. Journal of the Royal Statistical Society. Series D (The Statistician), 50 (2):218, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2681098.

Knight:2001:BRB

[164] John Knight. Book review: The Politics of Large Numbers: A History of Statistical Reasoning, by A. Desrosières; C. Naish. Journal of the Royal Statistical Society. Series D (The Statistician), 50 (2):218–219, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681099.

Kemp:2001:BRBe

[165] Freda Kemp. Book review: Essentials of Stochastic Processes, by R. Durrett. Journal of the Royal Statistical Society. Series D (The Statistician), 50 (2):220, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681100.

Wang:2001:BRB

[166] Duolao Wang. Book review: Nonparametric Curve Estimation: Methods, Theory, and Applications, by S. Efromovich. Journal of the Royal Statistical Society. Series D (The Statistician), 50(2):220–221, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681101.

Espejo:2001:BRBa

[167] Mariano Ruiz Espejo. Book review:

Guide to Economic Indicators, by N. Frumkin. Journal of the Royal Statistical Society. Series D (The Statistician), 50(2):221, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681102.

Kleiber:2001:BRB

[168] Christian Kleiber. Book review: Relative Distribution Methods in the Social Sciences, by M. S. Handcock; M. Morris. Journal of the Royal Statistical Society. Series D (The Statistician), 50 (2):222–223, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2681103.

Kemp:2001:BRBf

[169] A. W. Kemp. Book review: Interpretations of Probability, by A. Khrennikov. Journal of the Royal Statistical Society. Series D (The Statistician), 50 (2):223–225, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2681104.

Dialsingh:2001:BRBa

[170] Isaac Dialsingh. Book review: Mathematical Statistics, by K. Knight. Journal of the Royal Statistical Society. Series D (The Statistician), 50(2):225, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2681105.

Kemp:2001:BRBg

[171] Freda Kemp. Book review: An Introduction to Geometrical Probability: Distributional Aspects with Applications, by A. M. Mathai. Journal of the Royal Statistical Society. Series D (The Statistician), 50(2):225–226, ???? 2001. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2681106.

Burrell:2001:BRBb

[172] Quentin Burrell. Book review: Statistics, by J. T. McClave; T. Sincich; Chance Encounters: A First Course in Data Analysis and Inference, by C. J. Wild; G. A. F. Seber. Journal of the Royal Statistical Society. Series D (The Statistician), 50(2):226–228, ???? 2001.

CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2681107.

Kumar:2001:BRB

[173] Kuldeep Kumar. Book review: Introductory Mathematical Economics, by A. H. Mouhammed. Journal of the Royal Statistical Society. Series D (The Statistician), 50(2):228, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2681108.

Baxter:2001:BRB

[174] Mike Baxter. Book review: Sampling in Archaeology, by C. Orton. Journal of the Royal Statistical Society. Series D (The Statistician), 50(2):228–229, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2681109.

McKellar:2001:BRB

[175] John McKellar. Book review: How to Think about Statistics, by J. L. Phillips. Journal of the Royal Statistical Society. Series D (The Statistician), 50(2): 229–230, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681110.

Walters:2001:BRBa

[176] Stephen Walters. Book review: A Handbook of Statistical Analyses Using Stata, by S. Rabe-Hesketh; B. Everitt. Journal of the Royal Statistical Society. Series D (The Statistician), 50(2):230, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2681111.

Militino:2001:BRB

[177] Ana F. Militino. Book review: Linear Models in Statistics, by A. C. Rencher. Journal of the Royal Statistical Society. Series D (The Statistician), 50(2): 230–231, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2681112.

Mowll:2001:BRBa

[178] Deborah Mowll. Book review: Statistical Methods for Quality Improvement, by T. P. Ryan. Journal of the Royal Statistical Society. Series D (The Statistician), 50(2):231–232, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2681113.

Martin:2001:BRB

[179] R. J. Martin. Book review: Spatial Statistics for Remote Sensing, by A. Stein; F. van der Meer; B. Gorte. Journal of the Royal Statistical Society. Series D (The Statistician), 50(2): 232–233, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2681114.

Woodward:2001:BRBb

[180] P. W. Woodward. Book review: Design and Analysis in Chemical Research, by R. L. Tranter. Journal of the Royal Statistical Society. Series D (The Statistician), 50(2):233–235, ???? 2001. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2681115.

Tighe:2001:BRB

[181] James Tighe. Book review: Drawing Inferences from Self-Selected Samples, by H. Wainer. Journal of the Royal Statistical Society. Series D (The Statistician), 50(2):235–236, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681116.

Smith:2001:BRB

[182] J. U. M. Smith. Book review: Model Building in Mathematical Programming, by H. P. Williams. Journal of the Royal Statistical Society. Series D (The Statistician), 50(2):236, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2681117.

Kemp:2001:BRBh

[183] A. W. Kemp. Book review: Thesaurus of Univariate Discrete Probability Distributions, by G. Wimmer; G. Altmann. Journal of the Royal Statistical Society. Series D (The Statistician), 50(2): 236–238, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681118.

Davis:2001:BRB

[184] T. P. Davis. Book review: Review of Experiments: Planning, Analysis, and Parameter Design Optimization, by C. F. J. Wu; M. Hamada. Journal of the Royal Statistical Society. Series D (The Statistician), 50(2):238–239, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2681119.

Kemp:2001:BRBi

[185] Freda Kemp. Book review: CRC Standard Probability and Statistics Tables and Formulae, by D. Zwillinger; S. Kokoska. Journal of the Royal Statistical Society. Series D (The Statistician), 50(2):239–241, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2681120.

Anonymous:2001:BMb

[186] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 50(2):??, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2681121.

Consonni:2001:MDD

[187] G. Consonni and V. Leucari. Model determination for directed acyclic graphs. Journal of the Royal Statistical Society. Series D (The Statistician), 50(3): 243–256, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2680929.

Tunaru:2001:MAV

[188] Radu Tunaru. Models of association versus causal models for contingency tables. Journal of the Royal Statistical Society. Series D (The Statistician), 50 (3):257–269, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2680930.

Tan:2001:DIO

[189] Frans E. S. Tan, Mario J. N. Ouwens, and Martijn P. F. Berger. Detection of influential observations in longitudinal mixed effects regression models. Journal of the Royal Statistical Society. Series D (The Statistician), 50(3): 271–284, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2680931.

Gudmundsson:2001:ECF

[190] Gudmundur Gudmundsson. Estimation of continuous flows from observed aggregates. Journal of the Royal Statistical Society. Series D (The Statistician), 50 (3):285–293, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2680932.

Penny:2001:CMO

[191] Kay I. Penny and Ian T. Jolliffe. A comparison of multivariate outlier detection methods for clinical laboratory safety data. Journal of the Royal Statistical Society. Series D (The Statistician), 50 (3):295–308, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2680933.

Basu:2001:ISS

[192] Srabashi Basu. Improved small sample inference procedures for epidemiological parameters under cross-sectional sampling. Journal of the Royal Statistical Society. Series D (The Statistician), 50(3):309–319, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2680934.

McAlevey:2001:EBS

[193] Lynn McAlevey, André M. Everett, and Charles Sullivan. Evolution in business statistics curricula: Learning from the ‘Making Statistics More Effective in Schools of Business’ Conference. Journal of the Royal Statistical Society. Series D (The Statistician), 50 (3):321–333, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2680935.

Hinkelmann:2001:POK

[194] Klaus Hinkelmann. Professor Oscar Kempthorne, 1919–2000. Journal of the Royal Statistical Society. Series D (The Statistician), 50(3):335–336, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2680936.

Atkinson:2001:DER

[195] Tony Atkinson. Dr Ernest Rudd, 1924– 1999. Journal of the Royal Statistical Society. Series D (The Statistician), 50 (3):336–337, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2680937.

Anonymous:2001:FMc

[196] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 50(3):339, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2680928.

Stoyanov:2001:BRBb

[197] Jordan Stoyanov. Book review: Random Data: Analysis and Measurement Procedures, by J. S. Bendat; A. G. Piersol. Journal of the Royal Statistical Society. Series D (The Statistician), 50 (3):340, ???? 2001. CODEN ????

ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2680938.

Crowder:2001:BRB

[198] Martin Crowder. Book review: Reliability: Modelling, Prediction, and Optimization, by W. R. Blischke; D. N. P. Murthy. Journal of the Royal Statistical Society. Series D (The Statistician), 50(3):340–341, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2680939.

Congdon:2001:BRB

[199] Peter Congdon. Book review: Computer-Assisted Analysis of Mixtures and Applications: Meta-Analysis, Disease Mapping and Others, by D. Böhning. Journal of the Royal Statistical Society. Series D (The Statistician), 50(3):341–342, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2680940.

Balding:2001:BRB

[200] David Balding. Book review: Computational Molecular Biology, by P. Clote; R. Backofen. Journal of the Royal Statistical Society. Series D (The Statistician), 50(3):342–343, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2680941.

Zimeras:2001:BRB

[201] Stelios Zimeras. Book review: Generalized Linear Models: A Bayesian Perspective, by D. K. Dey; S. K. Ghosh; B. K. Mallick. Journal of the Royal Statistical Society. Series D (The Statistician), 50(3):343–344, ???? 2001. CODEN ????

ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2680942.

Scott:2001:BRBb

[202] Marian Scott. Book review: Item Response Theory for Psychologists, by S. E. Embretson; S. P. Reise. Journal of the Royal Statistical Society. Series D (The Statistician), 50(3):344–345, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2680943.

Walters:2001:BRBb

[203] Stephen Walters. Book review: Quality of Life: Assessment, Analysis and Interpretation, by P. M. Fayers; D. Machin. Journal of the Royal Statistical Society. Series D (The Statistician), 50(3): 345–346, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2680944.

Espejo:2001:BRBb

[204] Mariano Ruiz Espejo. Book review: Research Methods in Applied Settings: An Integrated Approach to Design and Analysis, by J. A. Gliner; G. A. Morgan. Journal of the Royal Statistical Society. Series D (The Statistician), 50 (3):346–347, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2680945.

Clarke:2001:BRB

[205] G. M. Clarke. Book review: A Course in Categorical Data Analysis, by T. Leonard. Journal of the Royal Statistical Society. Series D (The Statistician), 50(3):347–348, ???? 2001. CODEN ????

ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2680946.

Kiesel:2001:BRB

[206] Rüdiger Kiesel. Book review: Local Regression and Likelihood, by C. Loader. Journal of the Royal Statistical Society. Series D (The Statistician), 50(3):348, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2680947.

Hewson:2001:BRBc

[207] Paul Hewson. Book review: Statistics in Management Sciences, by A. Metcalfe. Journal of the Royal Statistical Society. Series D (The Statistician), 50 (3):348–349, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2680948.

Hewson:2001:BRBd

[208] Paul Hewson. Book review: Research and Evaluation for Business, by M. K. Pelosi; T. M. Sandifer; U. Sekaran. Journal of the Royal Statistical Society. Series D (The Statistician), 50(3): 349–350, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2680949.

Huson:2001:BRBa

[209] Les Huson. Book review: Medical Statistics at a Glance, by A. Petrie; C. Sabin. Journal of the Royal Statistical Society. Series D (The Statistician), 50(3): 350–351, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2680950.

Wood:2001:BRB

[210] Simon Wood. Book review: Mixed-Effects Models in S and S-Plus, by J. C. Pinheiro; D. M. Bates. Journal of the Royal Statistical Society. Series D (The Statistician), 50(3):351–352, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2680951.

Kemp:2001:BRBj

[211] Freda Kemp. Book review: Applied Statistical Decision Theory, by H. Raiffa; R. Schlaifer; Understanding Robust and Exploratory Data Analysis, by D. C. Hoaglin; F. Mosteller; J. W. Tukey. Journal of the Royal Statistical Society. Series D (The Statistician), 50(3): 352–353, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2680952.

Scott:2001:BRBc

[212] Marian Scott. Book review: A First Course in Structural Equation Modelling, by T. Raykov; G. A. Marcoulides. Journal of the Royal Statistical Society. Series D (The Statistician), 50(3): 353–354, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2680953.

Ravi:2001:BRB

[213] Sreenivasan Ravi. Book review: Essential Statistics, by D. G. Rees. Journal of the Royal Statistical Society. Series D (The Statistician), 50(3):354–355, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2680954.

Karlis:2001:BRB

[214] Dimitris Karlis. Book review: Statistics for People Who (Think They) Hate Statistics, by N. J. Salkind. Journal of the Royal Statistical Society. Series D (The Statistician), 50(3):355–356, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2680955.

Gani:2001:BRBa

[215] Joe Gani. Book review: Introduction to Stochastic Networks, by R. Serfozo. Journal of the Royal Statistical Society. Series D (The Statistician), 50(3): 356–358, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2680956.

Mowll:2001:BRBb

[216] Deborah Mowll. Book review: Statistical Methods in Software Engineering: Reliability and Risk, by N. D. Singpurwalla; S. P. Wilson. Journal of the Royal Statistical Society. Series D (The Statistician), 50(3):358, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2680957.

Newell:2001:BRB

[217] J. Newell. Book review: Multilevel Analysis: An Introduction to Basic and Advanced Multilevel Modeling, by T. A. B. Snijders; R. J. Bosker. Journal of the Royal Statistical Society. Series D (The Statistician), 50(3):358–359, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2680958.

Kemp:2001:BRBk

[218] David Kemp. Book review: Simulation: A Modeler’s Approach, by J. R. Thompson. Journal of the Royal Statistical Society. Series D (The Statistician), 50 (3):359–360, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2680959.

Brix:2001:BRB

[219] Anders Brix. Book review: Coupling, Stationarity and Regeneration, by H. Thorisson. Journal of the Royal Statistical Society. Series D (The Statistician), 50(3):360–361, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2680960.

Kemp:2001:BRBl

[220] A. W. Kemp. Book review: Econometric Analysis of Count Data, by R. Winkelmann. Journal of the Royal Statistical Society. Series D (The Statistician), 50 (3):361–362, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2680961.

Vencatasawmy:2001:BRB

[221] Coomaren P. Vencatasawmy. Book review: Introduction to Time Series Analysis and Forecasting with Applications of SAS and SPSS, by R. Yaffee; M. McGee. Journal of the Royal Statistical Society. Series D (The Statistician), 50(3):362–363, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2680962.

Anonymous:2001:CML

[222] Anonymous. Corrigendum: The median lethal dose — design and estimation. Journal of the Royal Statistical Society. Series D (The Statistician), 50 (3):364–366, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2680963. See [125].

Anonymous:2001:BMc

[223] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 50(3):??, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2680964.

Lievesley:2001:MDR

[224] Denise Lievesley. Making a difference:

A role for the responsible international statistician? Journal of the Royal Statistical Society. Series D (The Statistician), 50(4):367–406, ???? 2001. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2681223.

Lipsitz:2001:ECS

[225] Stuart R. Lipsitz, Nan M. Laird, Troyen A. Brennan, and Michael Parzen. Estimating the κ-coefficient from a selected sample. Journal of the Royal Statistical Society. Series D (The Statistician), 50(4):407–416, ???? 2001. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2681224.

Woodward:2001:GAI

[226] Phil Woodward and Janice Branson.

A graphical approach for incorporating prior knowledge when determining a sample size for the assessment of batched products. Journal of the Royal Statistical Society. Series D (The Statistician), 50(4):417–426, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681225.

Cavanaugh:2001:DAI

[227] Joseph E. Cavanaugh and Jacob J. Oleson. A diagnostic for assessing the influence of cases on the prediction of missing data. Journal of the Royal Statistical Society. Series D (The Statistician), 50(4):427–440, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681226.

Holt:2001:ROS

[228] Tim Holt, Ian Diamond, and Marie Cruddas. Risk in official statistics: A case-study of the 2001 One-Number Census Project. Journal of the Royal Statistical Society. Series D (The Statistician), 50(4):441–456, ???? 2001. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2681227.

Rosenbaum:2001:PJR

[229] Sidney Rosenbaum. Precursors of the Journal of the Royal Statistical Society. Journal of the Royal Statistical Society. Series D (The Statistician), 50 (4):457–466, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681228.

Anonymous:2001:RCS

[230] Anonymous. Report of the Council for the session 2000–2001. Journal of the Royal Statistical Society. Series D (The Statistician), 50(4):467–544, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2681229.

Anonymous:2001:FMd

[231] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 50(4):545, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2681222.

Mitchell:2001:BRBb

[232] Lada Mitchell. Book review: Statistics for Social and Health Research with a Guide to SPSS, by G. Argyrous. Journal of the Royal Statistical Society. Series D (The Statistician), 50(4): 546–547, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681230.

Gerontidis:2001:BRB

[233] Ioannis I. Gerontidis. Book review: Linear Algebra and Linear Models, by R. B. Bapat. Journal of the Royal Statistical Society. Series D (The Statistician), 50 (4):547–548, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681231.

Huson:2001:BRBb

[234] Les Huson. Book review: Introduction to Medical Statistics, by M. Bland. Journal of the Royal Statistical Society. Series D (The Statistician), 50(4):548, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2681232.

Boshnakov:2001:BRBb

[235] Georgi Boshnakov. Book review: Fourier Analysis of Time Series: An Introduction, by P. Bloomfield. Journal of the Royal Statistical Society. Series D (The Statistician), 50(4):549, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2681233.

Lawson:2001:BRB

[236] Andrew B. Lawson. Book review: Spatial Epidemiology: Methods and Applications, by P. Elliott; J. C. Wakefield; N. G. Best; D. J. Briggs. Journal of the Royal Statistical Society. Series D (The Statistician), 50(4):549–551, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2681234.

Bah:2001:BRB

[237] Sulaiman Bah. Book review: Discovering Statistics Using SPSS for Windows: Advanced Techniques for Beginners, by A. Field. Journal of the Royal Statistical Society. Series D (The Statistician), 50(4):551–552, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2681235.

Northrop:2001:BRB

[238] Paul Northrop. Book review: Generalized Linear Models: A Unified Approach, by J. Gill. Journal of the Royal Statistical Society. Series D (The Statistician), 50(4):552–553, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2681236.

Kemp:2001:BRBm

[239] Freda Kemp. Book review: Inequalities in Health: The Evidence, by D. Gordon; M. Shaw; D. Dorling; G. Davey-Smith. Journal of the Royal Statistical Society. Series D (The Statistician), 50(4):553, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681237.

Gani:2001:BRBb

[240] Joe Gani. Book review: Markov Point Processes and Their Applications, by M. N. M. van Lieshout. Journal of the Royal Statistical Society. Series D (The Statistician), 50(4):553–554, ???? 2001. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2681238.

Contoyannis:2001:BRB

[241] Paul Contoyannis. Book review:

Simulation-Based Inference in Econometrics: Methods and Applications, by R. Mariano; T. Schuerman; M. J. Weeks. Journal of the Royal Statistical Society. Series D (The Statistician), 50(4): 554–555, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681239.

Forchini:2001:BRB

[242] Giovanni Forchini and Patrick Marsh.

Book review: Applications of Differential Geometry to Econometrics, by P. Marriott; M. Salmon. Journal of the Royal Statistical Society. Series D (The Statistician), 50(4):555–556, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2681240.

Dialsingh:2001:BRBb

[243] Isaac Dialsingh. Book review: Applying Regression and Correlation, by J. Miles; M. Shevlin. Journal of the Royal Statistical Society. Series D (The Statistician), 50(4):556, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884

                                        (electronic). URL http://www.jstor.
                                        org/stable/2681241.

Leonard:2001:BRB

[244] Thomas Leonard. Book review: Statlabs: Mathematical Statistics through Applications, by D. Nolan; O. Speed. Journal of the Royal Statistical Society. Series D (The Statistician), 50(4): 556–557, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2681242.

Kemp:2001:BRBn

[245] Freda Kemp. Book review: Tackling Inequalities: Where Are We Now and What Can Be Done?, by C. Pantazis; D. Gordon. Journal of the Royal Statistical Society. Series D (The Statistician), 50(4):557, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884

                                        (electronic). URL http://www.jstor.
                                        org/stable/2681243.

Cervino:2001:BRB

[246] Alessandra Cervino. Book review: Practical Strategies for Experimenting, by G. K. Robinson. Journal of the Royal Statistical Society. Series D (The Statistician), 50(4):557–558, ???? 2001. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2681244.

Stoyanov:2001:BRBc

[247] Jordan Stoyanov. Book review: Topics in Finite and Discrete Mathematics, by S. M. Ross. Journal of the Royal Statistical Society. Series D (The Statistician), 50(4):558, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2681245.

Auton:2001:BRB

[248] Tim Auton. Book review: Modelling Survival Data: Extending the Cox Model, by T. M. Therneau; P. M. Grambsch. Journal of the Royal Statistical Society. Series D (The Statistician), 50(4):558–559, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681246.

Larsen:2001:BRB

[249] Pia Veldt Larsen. Book review: Asymptotics Statistics, by A. W. van der Vaart. Journal of the Royal Statistical Society. Series D (The Statistician), 50(4): 559–560, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681247.

Anonymous:2001:CBR

[250] Anonymous. Correction: Book reviews. Journal of the Royal Statistical Society. Series D (The Statistician), 50(4):560, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2681249.

Stoyanov:2001:BRBd

[251] Jordan Stoyanov. Book review: Probability via Expectation, by P. Whittle. Journal of the Royal Statistical Society. Series D (The Statistician), 50 (4):560, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2681248.

Anonymous:2001:VI

[252] Anonymous. Volume information. Journal of the Royal Statistical Society. Series D (The Statistician), 50(4):561–565, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2681221.

Anonymous:2001:BMd

[253] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 50(4):??, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2681250.

Anonymous:2002:FMa

[254] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 51(1):i–ii, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/3650385.

Chatfield:2002:CPS

[255] Chris Chatfield. Confessions of a pragmatic statistician. Journal of the Royal Statistical Society. Series D (The Statistician), 51(1):1–20, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884

                                        (electronic). URL http://www.jstor.
                                        org/stable/3650386.

Lim:2002:CTS

[256] Dong Hoon Lim and Seung Ju Jang. Comparison of two-sample tests for edge detection in noisy images. Journal of the Royal Statistical Society. Series D (The Statistician), 51(1):21–30, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/3650387.

Laverty:2002:SHM

[257] W. H. Laverty, M. J. Miket, and I. W.

Kelly. Simulation of hidden Markov models with EXCEL. Journal of the Royal Statistical Society. Series D (The Statistician), 51(1):31–40, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/3650388.

Jones:2002:SSD

[258] M. C. Jones. Student’s simplest distribution. Journal of the Royal Statistical Society. Series D (The Statistician), 51(1):41–49, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/3650389.

Franconi:2002:MBM

[259] Luisa Franconi and Julian Stander. A model-based method for disclosure limitation of business microdata. Journal of the Royal Statistical Society. Series D (The Statistician), 51(1):51–61, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
3650390.

vanDorp:2002:NET

[260] J. René van Dorp and Samuel Kotz.

A novel extension of the triangular distribution and its parameter estimation. Journal of the Royal Statistical Society. Series D (The Statistician), 51(1):

63–79, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/3650391.

Oakley:2002:EGP

[261] Jeremy Oakley. Eliciting Gaussian process priors for complex computer codes. Journal of the Royal Statistical Society. Series D (The Statistician), 51(1): 81–97, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/3650392.

Wood:2002:NAT

[262] Graham R. Wood and David J. Saville. A new angle on the t-test. Journal of the Royal Statistical Society. Series D (The Statistician), 51(1):99–104, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/3650393.

Barabesi:2002:NEC

[263] Lucio Barabesi and Luigi Greco. A note on the exact computation of the Student t, Snedecor F and sample correlation coefficient distribution functions. Journal of the Royal Statistical Society. Series D (The Statistician), 51(1): 105–110, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/3650394.

Firth:2002:OMG

[264] Catherine Firth and Wulf Rudoe. Obituary: Muriel Gore Venning, 1921–1997. Journal of the Royal Statistical Society. Series D (The Statistician), 51(1): 111–112, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/3650395.

Anonymous:2002:BMa

[265] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 51(1):113–126, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
3650410.

Burrell:2002:BRBa

[266] Quentin Burrell. Book review: Bayesian Theory, by J. M. Bernardo; A. F. M. Smith. Journal of the Royal Statistical Society. Series D (The Statistician), 51(1):114, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/3650396.

Edwards:2002:BRB

[267] A. W. F. Edwards. Book review:

The Mathematical Theory of Selection, Recombination, and Mutation, by R. Bürger. Journal of the Royal Statistical Society. Series D (The Statistician), 51(1):114–116, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/3650397.

Auton:2002:BRBa

[268] Tim Auton. Book review: The Theory of the Design of Experiments, by D. R. Cox; N. Reid. Journal of the Royal Statistical Society. Series D (The Statistician), 51(1):116, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/3650398.

Wooff:2002:BRB

[269] David Wooff. Book review: Introduction to Graphical Modelling, by D. Edwards. Journal of the Royal Statistical Society. Series D (The Statistician), 51 (1):116–117, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/3650399.

Knight:2002:BRBa

[270] John Knight. Book review: Foreign Direct Investment in Russia: A Strategy for Industrial Recovery, by P. Fischer. Journal of the Royal Statistical Society. Series D (The Statistician), 51 (1):117–118, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/3650400.

Dassios:2002:BRB

[271] Angelos Dassios. Book review: Derivatives in Financial Markets with Stochastic Volatility, by J.-P. Fouque; G. Papanicolaou; K. R. Sircar. Journal of the Royal Statistical Society. Series D (The Statistician), 51(1):118–119, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/3650401.

Sreenivasan:2002:BRBa

[272] Ravi Sreenivasan. Book review: Multiple and Generalized Nonparametric Regression, by J. Fox. Journal of the Royal Statistical Society. Series D (The Statistician), 51(1):119–120, ???? 2002. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/3650402.

Haigh:2002:BRBa

[273] John Haigh. Book review: Statistical Science in the Courtroom, by J. L. Gastwirth. Journal of the Royal Statistical Society. Series D (The Statistician), 51 (1):120–121, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/3650403.

Kemp:2002:BRBa

[274] Freda Kemp. Book review: Taking Chances: Winning with Probability, by J. Haigh. Journal of the Royal Statistical Society. Series D (The Statistician), 51(1):121, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/3650404.

Kemp:2002:BRBb

[275] Freda Kemp. Book review: The Job-Generation Controversy: The Economic Myth of Small Business, by D. Hirschberg. Journal of the Royal Statistical Society. Series D (The Statistician), 51(1):121–122, ???? 2002. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/3650405.

Downie:2002:BRB

[276] Tim Downie. Book review: Wavelet Methods for Time Series Analysis, by D. P. Percival; A. P. Walden. Journal of the Royal Statistical Society. Series D (The Statistician), 51(1):122–123, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/3650406.

Auton:2002:BRBb

[277] Tim Auton. Book review: The Analysis of Variance: Fixed, Random and Mixed Models, by H. Sahai; M. I. Ageel. Journal of the Royal Statistical Society. Series D (The Statistician), 51(1): 123–124, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/3650407.

Kumar:2002:BRBa

[278] Kuldeep Kumar. Book review: Time Series Analysis and Its Applications, by R. H. Shumway; D. S. Stoffer. Journal of the Royal Statistical Society. Series D (The Statistician), 51(1):124, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/3650408.

Stoimenova:2002:BRBa

[279] Eugenia Stoimenova. Book review: Computerized Adaptive Testing: A Primer, by H. Wainer. Journal of the Royal Statistical Society. Series D (The Statistician), 51(1):124–125, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/3650409.

Anonymous:2002:FMb

[280] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 51(2):i–ii, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/3650312.

Watkins:2002:E

[281] Alan Watkins and Linda Wolstenholme. Editorial. Journal of the Royal Statistical Society. Series D (The Statistician), 51(2):127–128, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/3650313.

Lee:2002:DLL

[282] Herbert K. H. Lee, Daniel L. Cork, and David J. Algranati. Did Lennox Lewis beat Evander Holyfield?: Methods for analysing small sample interrater agreement problems. Journal of the Royal Statistical Society. Series D (The Statistician), 51(2):129–146, ???? 2002. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/3650314.

Morton:2002:WWS

[283] R. Hugh Morton. Who won the Sydney 2000 Olympics?: An allometric approach. Journal of the Royal Statistical Society. Series D (The Statistician), 51 (2):147–155, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/3650315.

Crowder:2002:DMP

[284] Martin Crowder, Mark Dixon, Anthony Ledford, and Mike Robinson. Dynamic modelling and prediction of English football league matches for betting. Journal of the Royal Statistical Society. Series D (The Statistician), 51 (2):157–168, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/3650316.

Marchand:2002:CBS

[285] Éric Marchand. On the comparison between standard and random knockout tournaments. Journal of the Royal Statistical Society. Series D (The Statistician), 51(2):169–178, ???? 2002. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/3650317.

Cox:2002:ADD

[286] Trevor F. Cox and Ryan T. Dunn. An analysis of decathlon data. Journal of the Royal Statistical Society. Series D (The Statistician), 51(2):179–187, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
3650318.

Preston:2002:RRL

[287] Ian Preston and Jonathan Thomas. Rain rules for limited overs Cricket and probabilities of victory. Journal of the Royal Statistical Society. Series D (The Statistician), 51(2):189–202, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/3650319.

Haigh:2002:OSC

[288] John Haigh. Optimal strategy in casino stud poker. Journal of the Royal Statistical Society. Series D (The Statistician), 51(2):203–213, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/3650320.

Ketzscher:2002:EAE

[289] Robert Ketzscher and Trevor J. Ringrose. Exploratory analysis of European Professional Golf Association statistics. Journal of the Royal Statistical Society. Series D (The Statistician), 51(2):215–228, ???? 2002. CODEN ????

ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/3650321.

Forrest:2002:OUA

[290] David Forrest and Robert Simmons.

Outcome uncertainty and attendance demand in sport: The case of English soccer. Journal of the Royal Statistical Society. Series D (The Statistician), 51 (2):229–241, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/3650322.

Genest:2002:L

[291] Christian Genest, Richard A. Lockhart, and Michael A. Stephens. χ^{2} and the lottery. Journal of the Royal Statistical Society. Series D (The Statistician), 51 (2):243–257, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/3650323.

Eggar:2002:VFF

[292] M. H. Eggar. Validity of fitting a first-order Markov chain model to data. Journal of the Royal Statistical Society. Series D (The Statistician), 51(2): 259–265, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/3650324.

Avery:2002:FIM

[293] Peter J. Avery. Fitting interconnected Markov chain models: DNA sequences and test Cricket matches. Journal of the Royal Statistical Society. Series D (The Statistician), 51(2):267–278, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/3650325.

Hoaglin:2002:PJW

[294] David C. Hoaglin. Professor John W. Tukey, 1915–2000. Journal of the Royal Statistical Society. Series D (The Statistician), 51(2):279–280, ???? 2002. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/3650326.

Davis:2002:JEG

[295] James A. Davis. John Edward George Utting, 1916–2001. Journal of the Royal Statistical Society. Series D (The Statistician), 51(2):280–281, ???? 2002. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/3650327.

Anonymous:2002:BMb

[296] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 51(2):283–304, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        3650348.

Kelly:2002:BRB

[297] Gabrielle E. Kelly. Book review: Robust Diagnostic Regression Analysis, by A. Atkinson; M. Riani. Journal of the Royal Statistical Society. Series D (The Statistician), 51(2):284–285, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/3650328.

Coolen:2002:BRBa

[298] Frank Coolen. Book review: Decision Theory: An Introduction to Dynamic Programming and Sequential Decisions, by J. Bather. Journal of the Royal Statistical Society. Series D (The Statistician), 51(2):285–286, ???? 2002. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/3650329.

Barber:2002:BRBa

[299] Stuart Barber. Book review: Time-Series Forecasting, by C. Chatfield. Journal of the Royal Statistical Society. Series D (The Statistician), 51 (2):286, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/3650330.

Zimeras:2002:BRB

[300] Stelios Zimeras. Book review: The Theory of the Design of Experiments, by D. R. Cox; N. Reid. Journal of the Royal Statistical Society. Series D (The Statistician), 51(2):286–287, ???? 2002. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/3650331.

Kemp:2002:BRBc

[301] A. W. Kemp. Book review: Annotated Readings in the History of Statistics, by H. A. David; A. W. F. Edwards. Journal of the Royal Statistical Society. Series D (The Statistician), 51 (2):287–288, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/3650332.

Kemp:2002:BRBd

[302] David Kemp. Book review: Probability: Elements of the Mathematical Theory, by C. R. Heathcote. Journal of the Royal Statistical Society. Series D (The Statistician), 51(2):288–289, ???? 2002.

CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/3650333.

Thomas:2002:BRBa

[303] Ray Thomas. Book review: The Job-Generation Controversy: The Economic Myth of Small Business, by D. Hirschberg. Journal of the Royal Statistical Society. Series D (The Statistician), 51(2):289–291, ???? 2002. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/3650334.

Bartkowiak:2002:BRB

[304] Anna Bartkowiak. Book review: Econometric Modelling, Techniques and Applications, by S. Holly; M. Weale. Journal of the Royal Statistical Society. Series D (The Statistician), 51(2): 291–292, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/3650335.

Mitchell:2002:BRB

[305] Lada Mitchell. Book review: Applied Logistic Regression, by D. W. Hosmer; S. Lemeshow; Logistic Regression: A Primer, by F. C. Pampel. Journal of the Royal Statistical Society. Series D (The Statistician), 51(2):292–293, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/3650336.

Kiesel:2002:BRB

[306] Rüdiger Kiesel. Book review: Financial Derivatives in Theory and Practice, by P. J. Hunt; J. E. Kennedy. Journal of the Royal Statistical Society. Series D (The Statistician), 51(2):294, ???? 2002.

CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/3650337.

Kemp:2002:BRBe

[307] Adrienne W. Kemp. Book review: Extreme Value Distributions: Theory and Applications, by S. Kotz; S. Nadarajah; An Introduction to Statistical Modeling of Extreme Values, by S. Coles. Journal of the Royal Statistical Society. Series D (The Statistician), 51(2): 294–296, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/3650338.

Barber:2002:BRBb

[308] Stuart Barber. Book review: The Basics of S and S-PLUS, by A. Krause; M. Olson. Journal of the Royal Statistical Society. Series D (The Statistician), 51 (2):296–297, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/3650339.

Vencatasawmy:2002:BRB

[309] Coomaren P. Vencatasawmy. Book review: Principles of Multivariate Analysis: A User’s Perspective, by W. J. Krzanowski. Journal of the Royal Statistical Society. Series D (The Statistician), 51(2):297, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/3650340.

Larsen:2002:BRBa

[310] Pia Veldt Larsen. Book review: Asymptotics in Statistics, Some Basic Concepts, by L. Le Cam; G. L. Yang. Journal of the Royal Statistical Society. Series D (The Statistician), 51(2):

297–298, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/3650341.

Espejo:2002:BRBa

[311] Mariano Ruiz Espejo. Book review: Fluctuations in Physical Systems, by H. L. Pécseli. Journal of the Royal Statistical Society. Series D (The Statistician), 51(2):298–299, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/3650342.

Moreno:2002:BRB

[312] Victor Moreno. Book review: A User’s Guide to MLwiN (Version 2.1), by J. Rasbash; W. Browne; H. Goldstein; M. Yang; I. Plewis; M. Healy; G. Woodhouse; D. Draper; I. Langford; T. Lewis. Journal of the Royal Statistical Society. Series D (The Statistician), 51(2): 299–300, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/3650343.

Burrell:2002:BRBb

[313] Quentin Burrell. Book review: Wavelets: The Key to Intermittent Information?, by B. W. Silverman; J. C. Vassilicos. Journal of the Royal Statistical Society. Series D (The Statistician), 51(2):300–301, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/3650344.

Youngman:2002:BRB

[314] James H. R. Youngman. Book review: Basic Engineering Data Collection and Analysis, by S. B. Vardeman; J. M.

Jobe. Journal of the Royal Statistical Society. Series D (The Statistician), 51 (2):301–302, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/3650345.

Fernandez:2002:BRB

[315] Carmen Fernández. Book review: Modern Applied Statistics with S-PLUS, by W. N. Venables; B. D. Ripley. Journal of the Royal Statistical Society. Series D (The Statistician), 51(2): 302–303, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/3650346.

Thomas:2002:BRBb

[316] L. Thomas. Book review: S Programming, by W. N. Venables; B. D. Ripley. Journal of the Royal Statistical Society. Series D (The Statistician), 51 (2):303, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/3650347.

Anonymous:2002:FMc

[317] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 51(3):i–ii, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/3650275.

Langsrud:2002:MAV

[318] Øyvind Langsrud. 50–50 multivariate analysis of variance for collinear responses. Journal of the Royal Statistical Society. Series D (The Statistician), 51(3):305–317, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/3650276.

Lynn:2002:SRB

[319] Peter Lynn and Paul Clarke. Separating refusal bias and non-contact bias: Evidence from UK national surveys. Journal of the Royal Statistical Society. Series D (The Statistician), 51(3): 319–333, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/3650277.

Cinca:2002:MSE

[320] C. Serrano Cinca, C. Mar Molinero, and J. L. Gallizo Larraz. A multivariate study of the economy of the European Union via financial statements analysis. Journal of the Royal Statistical Society. Series D (The Statistician), 51 (3):335–354, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/3650278.

Patefield:2002:FNL

[321] Mike Patefield. Fitting non-linear structural relationships using SAS procedure NLMIXED. Journal of the Royal Statistical Society. Series D (The Statistician), 51(3):355–366, ???? 2002. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/3650279.

Donner:2002:TED

[322] Allan Donner and Guangyong Zou. Testing the equality of dependent intraclass correlation coefficients. Journal of the Royal Statistical Society. Series D (The Statistician), 51(3):367–379, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/3650280.

Basu:2002:IPM

[323] Ayanendranath Basu, Surajit Ray, Chanseok Park, and Srabashi Basu. Improved power in multinomial goodnessof-fit tests. Journal of the Royal Statistical Society. Series D (The Statistician), 51(3):381–393, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/3650281.

Caroni:2002:C

[324] C. Caroni, P. Prescott, Ian Jolliffe, and Kay Penny. Correspondence. Journal of the Royal Statistical Society. Series D (The Statistician), 51(3):395–397, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
3650282.

Gani:2002:PMB

[325] Joe Gani. Professor M. S. Bartlett FRS, 1910–2002. Journal of the Royal Statistical Society. Series D (The Statistician), 51(3):399–402, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/3650283.

Freeman:2002:CS

[326] Christopher Freeman. Christopher Saunders, 1907–2000. Journal of the Royal Statistical Society. Series D (The Statistician), 51(3):402–403, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/3650284.

Edwards:2002:PCB

[327] A. W. F. Edwards. Professor C. A. B. Smith, 1917–2002. Journal of the Royal Statistical Society. Series D (The Statistician), 51(3):404–405, ???? 2002. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/3650285.

Anonymous:2002:BMc

[328] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 51(3):407, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/3650311.

Kemp:2002:BRBf

[329] Adrienne W. Kemp. Book review: The Moment of Proof: Mathematical Epiphanies, by D. C. Benson. Journal of the Royal Statistical Society. Series D (The Statistician), 51(3):408, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/3650286.

Shoesmith:2002:BRBa

[330] Eddie Shoesmith. Book review: Structural Equation Modeling with AMOS: Basic Concepts, Applications, and Programming, by B. M. Byrne. Journal of the Royal Statistical Society. Series D (The Statistician), 51(3):408–409, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        3650287.

Tighe:2002:BRB

[331] James Tighe. Book review: Statistical Consulting: A Guide to Effective Communication, by J. Derr. Journal of the Royal Statistical Society. Series D (The Statistician), 51(3):409–410, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/3650288.

Jones:2002:BRBa

[332] M. C. Jones. Book review: Combinatorial Methods in Density Estimation, by L. Devroye; G. Lugosi. Journal of the Royal Statistical Society. Series D (The Statistician), 51(3):410–411, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/3650289.

Tolvi:2002:BRB

[333] Jussi Tolvi. Book review: Mathematics for Econometrics, by P. J. Dhrymes. Journal of the Royal Statistical Society. Series D (The Statistician), 51(3): 411–412, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/3650290.

Shoesmith:2002:BRBb

[334] Eddie Shoesmith. Book review: Introducing LISREL: A Guide for the Uninitiated, by A. Diamantopoulos; J. A. Siguaw. Journal of the Royal Statistical Society. Series D (The Statistician), 51(3):412–413, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/3650291.

Hemmett:2002:BRB

[335] Lindsay Hemmett. Book review: Beginning Statistics: An Introduction for Social Scientists, by I. Diamond; J. Jefferies. Journal of the Royal Statistical Society. Series D (The Statistician), 51 (3):413–414, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/3650292.

Jones:2002:BRBb

[336] M. C. Jones. Book review: Correlation and Dependence, by D. Drouet Mari; S. Kotz. Journal of the Royal Statistical Society. Series D (The Statistician), 51 (3):414–415, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/3650293.

Espejo:2002:BRBb

[337] Mariano Ruiz Espejo. Book review: Learning and Expectations in Macroeconomics, by G. W. Evans; S. Honkapohja. Journal of the Royal Statistical Society. Series D (The Statistician), 51 (3):415–416, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/3650294.

Sreenivasan:2002:BRBb

[338] Ravi Sreenivasan. Book review: Applications of Empirical Process Theory, by S. A. van de Geer. Journal of the Royal Statistical Society. Series D (The Statistician), 51(3):416–417, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/3650295.

Knight:2002:BRBb

[339] John Knight. Book review: Breadline Europe: The Measurement of Poverty, by D. Gordon; P. Townsend. Journal of the Royal Statistical Society. Series D (The Statistician), 51(3):417–418, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
3650296.

Bolton:2002:BRB

[340] Richard Bolton. Book review:

XploRe(R): Application Guide, by W. Härdle; Z. Hlávka; S. Klinke. Journal of the Royal Statistical Society. Series D (The Statistician), 51(3):418–419, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
3650297.

Hilliam:2002:BRB

[341] Rachel Hilliam. Book review: Time Reversibility, Computer Simulation, and Chaos, by W. G. Hoover. Journal of the Royal Statistical Society. Series D (The Statistician), 51(3):419, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/3650298.

Espejo:2002:BRBc

[342] Mariano Ruiz Espejo. Book review:

Essential Data Skills for Business and Management, by S. Hurd; J. Mangan. Journal of the Royal Statistical Society. Series D (The Statistician), 51(3): 419–421, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/3650299.

Kemp:2002:BRBg

[343] Malcolm Kemp. Book review: Option Pricing, Interest Rates and Risk Management, by E. Jouini; J. Cvitani; M. Musiela. Journal of the Royal Statistical Society. Series D (The Statistician), 51(3):421, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/3650300.

Bartolucci:2002:BRB

[344] Francesco Bartolucci and Nicola Falocci. Book review: A Guide to Monte Carlo Simulations in Statistical Physics, by D. P. Landau; K. Binder. Journal of the Royal Statistical Society. Series D (The Statistician), 51(3):422, ???? 2002. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/3650301.

Molchanov:2002:BRBa

[345] Ilya Molchanov. Book review: An Invariant Approach to Statistical Analysis of Shapes, by S. R. Lele; J. T. Richtsmeier. Journal of the Royal Statistical Society. Series D (The Statistician), 51(3):422–423, ???? 2002. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/3650302.

Molchanov:2002:BRBb

[346] Ilya Molchanov. Book review: Statistics of Random Processes, Vols I, II, by R. S. Liptser; A. N. Shiryaev. Journal of the Royal Statistical Society. Series D (The Statistician), 51(3):423, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/3650303.

Kemp:2002:BRBh

[347] Adrienne W. Kemp. Book review: Finite Mixture Models, by G. J. McLachlan; D. Peel. Journal of the Royal Statistical Society. Series D (The Statistician), 51(3):423–424, ???? 2002. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/3650304.

Jones:2002:BRBc

[348] M. C. Jones. Book review: Permutation Methods: A Distance Function Approach, by P. W. Mielke; K. J. Berry. Journal of the Royal Statistical Society. Series D (The Statistician), 51(3):425, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/3650305.

Lawson:2002:BRB

[349] Andrew Lawson. Book review: Environmental Statistics with S-Plus, by S. P. Millard; N. K. Neerchal. Journal of the Royal Statistical Society. Series D (The Statistician), 51(3):425–426, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/3650306.

Morris:2002:BRBa

[350] Clare Morris. Book review: SPSS for Windows: An Introduction to Use and Interpretation in Research, by G. A. Morgan; O. V. Griego; G. W. Gloeckner. Journal of the Royal Statistical Society. Series D (The Statistician), 51 (3):426–427, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/3650307.

Kemp:2002:BRBi

[351] Adrienne W. Kemp. Book review: Spatial Tessellations: Concepts and Applications of Voronoi Diagrams, by A. Okabe; B. Boots; K. Sugihara; S. N. Chui. Journal of the Royal Statistical Society. Series D (The Statistician), 51(3):

427–428, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/3650308.

Barber:2002:BRBc

[352] Stuart Barber. Book review: A Course in Time Series Analysis, by D. Peña; G. C. Tiao; R. S. Tsay. Journal of the Royal Statistical Society. Series D (The Statistician), 51(3):428–429, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/3650309.

Kemp:2002:BRBj

[353] Freda Kemp. Book review: Finite Population Sampling and Inference: A Prediction Approach, by R. Valliant; A. H. Dorfman; R. M. Royall. Journal of the Royal Statistical Society. Series D (The Statistician), 51(3):429–430, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/3650310.

Anonymous:2002:FMd

[354] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 51(4):i–ii, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/3650350.

Laws:2002:HBM

[355] David J. Laws and Anthony O’Hagan. A hierarchical Bayes model for multilocation auditing. Journal of the Royal Statistical Society. Series D (The Statistician), 51(4):431–450, ???? 2002. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/3650351.

DeSantis:2002:IUD

[356] Fulvio De Santis. Interactive use of default and robust Bayes testing methods in the presence of vague prior information. Journal of the Royal Statistical Society. Series D (The Statistician), 51 (4):451–465, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/3650352.

Klar:2002:EBC

[357] Neil Klar, Stuart R. Lipsitz, Michael Parzen, and Traci Leong. An exact bootstrap confidence interval for κinsmallsamples.JournaloftheRoyalStatisticalSociety.SeriesD(TheStatistician), 51 (4) : 467−−478,????2002.CODEN????ISSN0039− 0526(print), 1467−9884(electronic).URLhttpwwwjstororgstable3650353.

Ukoumunne:2002:NUV

[358] Obioha C. Ukoumunne, Martin C. Gulliford, and Susan Chinn. A note on the use of the variance inflation factor for determining sample size in cluster randomized trials. Journal of the Royal Statistical Society. Series D (The Statistician), 51(4):479–484, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/3650354.

Anonymous:2002:RCS

[359] Anonymous. Report of the Council for the session 2001–2002 Council report. Journal of the Royal Statistical Society. Series D (The Statistician), 51(4): 485–563, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/3650355.

Cleland:2002:PWB

[360] John Cleland. Professor William Brass, 1921–1999. Journal of the Royal Statistical Society. Series D (The Statistician), 51(4):565–566, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/3650356.

Armitage:2002:PHO

[361] Peter Armitage. Professor Henry Oliver Lancaster, 1913–2001. Journal of the Royal Statistical Society. Series D (The Statistician), 51(4):566–568, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/3650357.

Styan:2002:HR

[362] George P. H. Styan. Harold Ruben, 1923–2001. Journal of the Royal Statistical Society. Series D (The Statistician), 51(4):568–570, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/3650358.

Anonymous:2002:BMd

[363] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 51(4):571, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/3650384.

Sanford:2002:BRB

[364] E. Sanford. Book review: Making Sense of Data: A Self-Instruction Manual on the Interpretation of Epidemiological Data, by J. H. Abramson; Z. H. Abramson. Journal of the Royal Statistical Society. Series D (The Statistician), 51(4):572, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/3650359.

Bashir:2002:BRB

[365] Saghir Bashir. Book review: Missing Data, by P. D. Allison. Journal of the Royal Statistical Society. Series D (The Statistician), 51(4):572–573, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/3650360.

Hall:2002:BRB

[366] Peter Hall. Book review: Statistical Data Analysis Using Your Personal Computer, by I. H. Bernstein; N. A. Rowe. Journal of the Royal Statistical Society. Series D (The Statistician), 51 (4):573–574, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/3650361.

Kemp:2002:BRBk

[367] Freda Kemp. Book review: Statistical Questions in Evidence-Based Medicine, by M. Bland; J. Peacock. Journal of the Royal Statistical Society. Series D (The Statistician), 51(4):574, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/3650362.

Swires-Hennessy:2002:BRB

[368] Ed Swires-Hennessy. Book review: Correlation and Regression: Applications for Industrial Organizational Psychology and Management, by P. Bobko. Journal of the Royal Statistical Society. Series D (The Statistician), 51(4): 574–575, ???? 2002. CODEN ????

ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/3650363.

Hewson:2002:BRB

[369] Paul Hewson. Book review: Practical Bayesian Modelling, by P. Congdon. Journal of the Royal Statistical Society. Series D (The Statistician), 51 (4):575–576, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/3650364.

Molchanov:2002:BRBc

[370] Ilya Molchanov. Book review: Best Approximation in Inner Product Spaces, by F. Deutsch. Journal of the Royal Statistical Society. Series D (The Statistician), 51(4):576, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/3650365.

Gray:2002:BRB

[371] Alison Gray. Book review: Approximating Integrals via Monte Carlo and Deterministic Methods, by M. Evans; T. Swartz. Journal of the Royal Statistical Society. Series D (The Statistician), 51(4):576–577, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/3650366.

Haigh:2002:BRBb

[372] John Haigh. Book review: Statistics for Lawyers, by M. O. Finkelstein; B. Levin. Journal of the Royal Statistical Society. Series D (The Statistician), 51 (4):577–578, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/3650367.

Kemp:2002:BRBl

[373] Freda Kemp. Book review: Statistical Decision Theory, by S. French; D. Rios Insua. Journal of the Royal Statistical Society. Series D (The Statistician), 51 (4):578–579, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/3650368.

Liao:2002:BRB

[374] Tim Futing Liao. Book review: Sociology: Numbers, Narratives, and Integration of Research and Theory, by J. H. Goldthorpe. Journal of the Royal Statistical Society. Series D (The Statistician), 51(4):579–580, ???? 2002. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/3650369.

Reale:2002:BRB

[375] Marco Reale. Book review: Practical Time Series, by G. Janacek. Journal of the Royal Statistical Society. Series D (The Statistician), 51(4):580, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/3650370.

Kumar:2002:BRBb

[376] Kuldeep Kumar. Book review: Structural Equation Modelling: Foundations and Extensions, by D. Kaplan. Journal of the Royal Statistical Society. Series D (The Statistician), 51(4):581, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/3650371.

McCulloch:2002:BRB

[377] Andrew McCulloch. Book review: Statistical Concepts: Education and Behavioral Sciences, by R. G. Lomax. Journal of the Royal Statistical Society. Series D (The Statistician), 51(4): 581–582, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/3650372.

French:2002:BRB

[378] Simon French. Book review: Statistics for Environmental Science and Management, by B. F. J. Manly. Journal of the Royal Statistical Society. Series D (The Statistician), 51(4):582–583, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/3650373.

Kemp:2002:BRBm

[379] David Kemp. Book review: The Subjectivity of Scientists and the Bayesian Approach, by S. J. Press; J. M. Tanur. Journal of the Royal Statistical Society. Series D (The Statistician), 51 (4):583–584, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/3650374.

Auton:2002:BRBc

[380] Tim Auton. Book review: Introducing ANOVA and ANCOVA, a GLM Approach, by A. Rutherford. Journal of the Royal Statistical Society. Series D (The Statistician), 51(4):584, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/3650375.

Larsen:2002:BRBb

[381] Pia Veldt Larsen. Book review: Likelihood Methods in Statistics, by T. A. Severini. Journal of the Royal Statistical Society. Series D (The Statistician), 51(4):584–585, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/3650376.

Burrell:2002:BRBc

[382] Quentin Burrell. Book review: Calculated Bets: Computers, Gambling, and Mathematical Modeling to Win, by S. Skiena. Journal of the Royal Statistical Society. Series D (The Statistician), 51(4):585–586, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/3650377.

Morris:2002:BRBb

[383] Clare Morris. Book review: SAS Programming for Researchers and Social Scientists, by P. E. Spector. Journal of the Royal Statistical Society. Series D (The Statistician), 51(4):586, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/3650378.

Coolen:2002:BRBb

[384] F. P. A. Coolen. Book review: Causation, Prediction, and Search, by P. Spirtes; C. Glymour; R. Scheines. Journal of the Royal Statistical Society. Series D (The Statistician), 51(4):586–587, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
3650379.

Chandler:2002:BRB

[385] Richard Chandler. Book review: Statistical Analysis in Climate Research, by H. von Storch; F. W. Zwiers. Journal of the Royal Statistical Society. Series D (The Statistician), 51(4):587–588, ???? 2002.

CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/3650380.

Stoimenova:2002:BRBb

[386] Eugenia Stoimenova. Book review: Test Scoring, by D. Thissen; H. Wainer. Journal of the Royal Statistical Society. Series D (The Statistician), 51(4): 588–589, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/3650381.

Auton:2002:BRBd

[387] Tim Auton. Book review: Single-Case and Small-N Experimental Designs: A Practical Guide to Randomization Tests, by J. B. Todman; P. Dugard. Journal of the Royal Statistical Society. Series D (The Statistician), 51(4): 589–590, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/3650382.

Sisson:2002:BRB

[388] Scott Sisson. Book review: Fundamentals of Modern Statistical Methods: Substantially Improving Power and Accuracy, by R. R. Wilcox. Journal of the Royal Statistical Society. Series D (The Statistician), 51(4):590–591, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/3650383.

Anonymous:2002:VI

[389] Anonymous. Volume information. Journal of the Royal Statistical Society. Series D (The Statistician), 51(4):593–598, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
3650349.

Haigh:2003:ERJ

[390] John Haigh and Alan Watkins. Editorial: Restructuring the “Journal of the Royal Statistical Society”. Journal of the Royal Statistical Society. Series D (The Statistician), 52(1):1–2, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/4128165.

Wright:2003:MMR

[391] David E. Wright and Isabelle Bray.

A mixture model for rounded data. Journal of the Royal Statistical Society. Series D (The Statistician), 52 (1):3–13, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/4128166.

Benton:2003:IMU

[392] Denise Benton, K. Krishnamoorthy, and Thomas Mathew. Inferences in multivariate-univariate calibration problems. Journal of the Royal Statistical Society. Series D (The Statistician), 52 (1):15–39, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/4128167.

Mudholkar:2003:CWR

[393] Govind S. Mudholkar and Gregory E.

Wilding. On the conventional wisdom regarding two consistent tests of bivariate independence. Journal of the Royal Statistical Society. Series D (The Statistician), 52(1):41–57, ???? 2003. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/4128168.

Singh:2003:LRR

[394] Housila P. Singh and Mariano Ruiz Espejo. On linear regression and ratioproduct estimation of a finite population mean. Journal of the Royal Statistical Society. Series D (The Statistician), 52(1):59–67, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884

                                        (electronic). URL http://www.jstor.
                                        org/stable/4128169.

Hayes:2003:SNS

[395] Kevin Hayes and Tony Kinsella. Spurious and non-spurious power in performance criteria for tests of discordancy. Journal of the Royal Statistical Society. Series D (The Statistician), 52 (1):69–82, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/4128170.

Novikov:2003:RES

[396] Ilya Novikov. A remark on efficient simulations in SAS. Journal of the Royal Statistical Society. Series D (The Statistician), 52(1):83–86, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884

                                        (electronic). URL http://www.jstor.
                                        org/stable/4128171.

Paulino:2003:ARI

[397] Carlos Daniel Paulino and Paulo Soares. Analysis of rates in incomplete Poisson data. Journal of the Royal Statistical Society. Series D (The Statistician), 52 (1):87–99, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/4128172.

Deny:2003:AHJ

[398] Brian J. Deny. Arnold H. J. Baines, 1921–2001. Journal of the Royal Statistical Society. Series D (The Statistician), 52(1):101–102, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/4128173.

Dunsmuir:2003:RT

[399] William T. M. Dunsmuir, Sean P.

Meyn, and Gareth O. Roberts. Richard Tweedie, 1947–2001. Journal of the Royal Statistical Society. Series D (The Statistician), 52(1):102–104, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/4128174.

Anonymous:2003:BMa

[400] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 52(1):105, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/4128200.

Welty:2003:BRB

[401] Gordon Welty. Book review: Principles of Forecasting: A Handbook for Researchers and Practitioners, by J. S. Armstrong. Journal of the Royal Statistical Society. Series D (The Statistician), 52(1):106–107, ???? 2003. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/4128175.

Kemp:2003:BRBa

[402] Freda Kemp. Book review: Advanced Linear Modeling, by R. Christensen. Journal of the Royal Statistical Society. Series D (The Statistician), 52 (1):107, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/4128176.

Kemp:2003:BRBb

[403] A. W. Kemp. Book review: Statistics in Psychology: An Historical Perspective, by M. Cowles. Journal of the Royal Statistical Society. Series D (The Statistician), 52(1):107–109, ???? 2003. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/4128177.

Kemp:2003:BRBc

[404] Freda Kemp. Book review: Epidemic Modelling: An Introduction, by D. J. Daley; J. M. Gani. Journal of the Royal Statistical Society. Series D (The Statistician), 52(1):109–110, ???? 2003. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/4128178.

Huson:2003:BRBa

[405] Les Huson. Book review: Clinical Trials, by L. Duley; B. Farrell. Journal of the Royal Statistical Society. Series D (The Statistician), 52(1):110–111, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/4128179.

Kemp:2003:BRBd

[406] Freda Kemp. Book review: The Cambridge Dictionary of Statistics, by B. S. Everitt. Journal of the Royal Statistical Society. Series D (The Statistician), 52(1):111–112, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/4128180.

Brix:2003:BRBa

[407] Anders Brix. Book review: Nonlinear and Nonstationary Signal Processing, by W. J. Fitzgerald; R. L. Smith; A. T. Walden; P. C. Young. Journal of the Royal Statistical Society. Series D (The Statistician), 52(1):112–113, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/4128181.

Espejo:2003:BRB

[408] Mariano Ruiz Espejo. Book review: Statistical Decision Theory, by S. French; D. Rios Insua. Journal of the Royal Statistical Society. Series D (The Statistician), 52(1):113, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/4128182.

Tolvi:2003:BRBa

[409] Jussi Tolvi. Book review: Essays in Econometrics-Collected Papers of Clive W. J. Granger, by E. Ghysels; N. R. Swanson; M. W. Watson. Journal of the Royal Statistical Society. Series D (The Statistician), 52(1):113–114, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/4128183.

Mitchell:2003:BRBa

[410] Lada Mitchell. Book review: Regression Modeling Strategies, by F. E. Harrell, Jr. Journal of the Royal Statistical Society. Series D (The Statistician), 52 (1):114–115, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/4128184.

Kemp:2003:BRBe

[411] Freda Kemp. Book review: Statisticians of the Centuries, by C. C. Heyde; E. Seneta. Journal of the Royal Statistical Society. Series D (The Statistician), 52(1):115–116, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/4128185.

French:2003:BRB

[412] Simon French. Book review: Bayesian Networks and Decision Graphs, by F. V. Jensen. Journal of the Royal Statistical Society. Series D (The Statistician), 52 (1):116–117, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/4128186.

Tighe:2003:BRBa

[413] James Tighe. Book review: Regression Basics, by L. H. Kahane. Journal of the Royal Statistical Society. Series D (The Statistician), 52(1):117–118, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/4128187.

DSouza:2003:BRBa

[414] J. C. D’Souza. Book review: A First Course on Parametric Inference, by B. K. Kale. Journal of the Royal Statistical Society. Series D (The Statistician), 52(1):118–119, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/4128188.

Burrell:2003:BRB

[415] Quentin Burrell. Book review: Should We Risk It?: Exploring Environmental, Health, and Technological Problem Solving, by D. M. Kammen; D. M. Hassenzahl. Journal of the Royal Statistical Society. Series D (The Statistician), 52 (1):119–120, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/4128189.

Kemp:2003:BRBf

[416] Freda Kemp. Book review: Data Analysis, by Resampling: Concepts and Applications, by C. E. Lunneborg. Journal of the Royal Statistical Society. Series D (The Statistician), 52(1):120–121, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
4128190.

Kemp:2003:BRBg

[417] David Kemp. Book review: Counting:

The Art of Enumerative Combinatorics, by G. E. Martin. Journal of the Royal Statistical Society. Series D (The Statistician), 52(1):121–122, ???? 2003. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/4128191.

Hewson:2003:BRBa

[418] Paul Hewson. Book review: Geographic Information Systems for Transportation: Principles and Applications, by H. J. Miller; S.-L. Shaw. Journal of the Royal Statistical Society. Series D (The Statistician), 52(1):122–123, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/4128192.

Swires-Hennessy:2003:BRBa

[419] Ed Swires-Hennessy. Book review: How to Forecast: A Guide for Business, by J.

Morrell. Journal of the Royal Statistical Society. Series D (The Statistician), 52(1):123–124, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/4128193.

Molchanov:2003:BRB

[420] Ilya Molchanov. Book review: Advanced Mean Field Methods, by M. Opper; D. Saad. Journal of the Royal Statistical Society. Series D (The Statistician), 52(1):124, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/4128194.

Gentle:2003:BRB

[421] James E. Gentle. Book review: Numerical Recipes in C++, by W. H. Press; S. A. Teukolsky; W. T. Vetterling; B. P. Flannery. Journal of the Royal Statistical Society. Series D (The Statistician), 52(1):124–125, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/4128195.

Coolen:2003:BRBa

[422] Frank Coolen. Book review: The Bayesian Choice, by C. P. Robert. Journal of the Royal Statistical Society. Series D (The Statistician), 52(1): 125–126, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/4128196.

Kemp:2003:BRBh

[423] Freda Kemp. Book review: Smoothing and Regression: Approaches, Computation, and Application, by M. G. Schimek. Journal of the Royal Statistical Society. Series D (The Statistician), 52(1):126–127, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/4128197.

Sisson:2003:BRBa

[424] Scott Sisson. Book review: Practical Statistics for Environmental and Biological Scientists, by J. Townend. Journal of the Royal Statistical Society. Series D (The Statistician), 52(1): 127–128, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/4128198.

Tolvi:2003:BRBb

[425] Jussi Tolvi. Book review: Analysis of Financial Time Series, by R. S. Tsay. Journal of the Royal Statistical Society. Series D (The Statistician), 52(1): 128–129, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/4128199.

Anonymous:2003:FMa

[426] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 52(1):??, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/4128164.

Pham-Gia:2003:DES

[427] Thu Pham-Gia and Noyan Turkkan. Determination of exact sample sizes in the Bayesian estimation of the difference of two proportions. Journal of the Royal Statistical Society. Series D (The Statistician), 52(2):131–150, ???? 2003. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/4128241.

Qiu:2003:NMC

[428] Peihua Qiu and Douglas Hawkins. A nonparametric multivariate cumulative sum procedure for detecting shifts in all directions. Journal of the Royal Statistical Society. Series D (The Statistician), 52(2):151–164, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/4128242.

Candolo:2003:NMU

[429] C. Candolo, A. C. Davison, and C. G. B. Demétrio. A note on model uncertainty in linear regression. Journal of the Royal Statistical Society. Series D (The Statistician), 52(2):165–177, ???? 2003. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/4128243.

Faraggi:2003:ARO

[430] David Faraggi. Adjusting receiver operating characteristic curves and related indices for covariates. Journal of the Royal Statistical Society. Series D (The Statistician), 52(2):179–192, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/4128244.

Jackson:2003:MMM

[431] Christopher H. Jackson, Linda D. Sharples, Simon G. Thompson, Stephen W. Duffy, and Elisabeth Couto. Multistate Markov models for disease progression with classification error. Journal of the Royal Statistical Society. Series D (The Statistician), 52(2):193–209, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/4128245.

Mellenbergh:2003:CIR

[432] Gideon J. Mellenbergh, Herman J. Adèr, Davis Baird, Martijn P. F. Berger, John E. Cornell, Jacques A. P. Hagenaars, and Peter C. M. Molenaar. Conceptual issues of research methodology for the behavioural, life and social sciences. Journal of the Royal Statistical Society. Series D (The Statistician), 52 (2):211–218, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/4128246.

Haigh:2003:WL

[433] John Haigh. The weakest link. Journal of the Royal Statistical Society. Series D (The Statistician), 52(2):219–226, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
4128247.

Nadarajah:2003:C

[434] Saralees Nadarajah, J. R. van Dorp, and S. Kotz. Correspondence. Journal of the Royal Statistical Society. Series D (The Statistician), 52(2):227–230, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/4128248.

Lindley:2003:PGB

[435] Dennis V. Lindley. Professor George A. Barnard, 1915–2002. Journal of the Royal Statistical Society. Series D (The Statistician), 52(2):231–233, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/4128249.

Rasbash:2003:IL

[436] Jon Rasbash. Ian Langford, 1961–2002. Journal of the Royal Statistical Society. Series D (The Statistician), 52(2): 233–234, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/4128250.

Anonymous:2003:BMb

[437] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 52(2):235, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/4128276.

Haigh:2003:BRBa

[438] John Haigh. Book review: Probabilistic Risk Analysis: Foundations and Methods, by T. Bedford; R. Cooke. Journal of the Royal Statistical Society. Series D (The Statistician), 52(2):236, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/4128251.

Tom:2003:BRB

[439] Brian D. M. Tom. Book review: Techniques of Event History Modeling: New Approaches to Causal Analysis, by H.- P. Blossfeld; G. Rohwer. Journal of the Royal Statistical Society. Series D (The Statistician), 52(2):236–238, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/4128252.

Tighe:2003:BRBb

[440] James Tighe. Book review: Statistical Consulting, by J. Cabrera; A. Mc-Dougall. Journal of the Royal Statistical Society. Series D (The Statistician), 52(2):238, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/4128253.

Kemp:2003:BRBi

[441] Freda Kemp. Book review: A Course in Probability Theory, by K. L. Chung. Journal of the Royal Statistical Society. Series D (The Statistician), 52 (2):239, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/4128254.

Rigby:2003:BRBa

[442] Alan S. Rigby. Book review: Analysis of Longitudinal Data, by P. J. Diggle; P. J. Heagerty; K.-Y. Liang; S. L. Zeger. Journal of the Royal Statistical Society. Series D (The Statistician), 52 (2):239–240, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/4128255.

Best:2003:BRBa

[443] Mark A. Best. Book review: Research Methods for Postgraduates, by T. Greenfield. Journal of the Royal Statistical Society. Series D (The Statistician), 52 (2):240–241, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/4128256.

Kemp:2003:BRBj

[444] Freda Kemp. Book review: Matrices for Statistics, by M. J. R. Healy. Journal of the Royal Statistical Society. Series D (The Statistician), 52(2):241, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/4128257.

Kemp:2003:BRBk

[445] Freda Kemp. Book review: Nonlinear Statistical Modeling, by C. Hsiao; K. Morimune; J. L. Powell. Journal of the Royal Statistical Society. Series D (The Statistician), 52(2):241–242, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/4128258.

Mitchell:2003:BRBb

[446] Lada Mitchell. Book review: Generalized, Linear, and Mixed Models, by C. E. McCulloch; S. R. Searle. Journal of the Royal Statistical Society. Series D (The Statistician), 52(2):242–243, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/4128259.

Brayshaw:2003:BRB

[447] Nigel Brayshaw. Book review: Applied Statistics in the Pharmaceutical Industry, with Case Studies in S-Plus, by S. P. Millard; A. Krause. Journal of the Royal Statistical Society. Series D (The Statistician), 52(2):243–244, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/4128260.

Auton:2003:BRB

[448] Tim Auton. Book review: Applied Stochastic Modelling, by B. J. T. Morgan. Journal of the Royal Statistical Society. Series D (The Statistician), 52 (2):244, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/4128261.

Swires-Hennessy:2003:BRBb

[449] Ed Swires-Hennessy. Book review:

Sources of Non-Official UK Statistics, by D. Mort. Journal of the Royal Statistical Society. Series D (The Statistician), 52(2):244–245, ???? 2003. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/4128262.

McConway:2003:BRB

[450] Kevin McConway. Book review: Generalized Linear Models with Applications in Engineering and the Sciences, by R. H. Myers; D. C. Montgomery; G. G. Vining. Journal of the Royal Statistical Society. Series D (The Statistician), 52(2):245–246, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/4128263.

Sreenivasan:2003:BRB

[451] Ravi Sreenivasan. Book review: Modeling in Medical Decision Making: A Bayesian Approach, by G. Parmigiani. Journal of the Royal Statistical Society. Series D (The Statistician), 52 (2):246, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/4128264.

Huson:2003:BRBb

[452] L. W. Huson. Book review: Multivariate Permutation Tests: With Applications in Biostatistics, by F. Pesarin. Journal of the Royal Statistical Society. Series D (The Statistician), 52(2):247, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/4128265.

Stoyanov:2003:BRBa

[453] Jordan Stoyanov. Book review: A Guide to First-Passage Processes, by S. Redner. Journal of the Royal Statistical Society. Series D (The Statistician), 52 (2):247, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/4128266.

Larsen:2003:BRBa

[454] Pia Veldt Larsen. Book review: Approximation Theorems of Mathematical Statistics, by R. J. Serfling. Journal of the Royal Statistical Society. Series D (The Statistician), 52(2):247–248, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        4128267.

Kemp:2003:BRBl

[455] Freda Kemp. Book review: The Handbook of Parametric and Nonparametric Statistical Procedures, by D. Sheskin. Journal of the Royal Statistical Society. Series D (The Statistician), 52(2): 248–249, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/4128268.

Kemp:2003:BRBm

[456] Freda Kemp. Book review: Probability for Statisticians, by G. R. Shorack. Journal of the Royal Statistical Society. Series D (The Statistician), 52 (2):249, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/4128269.

Sprent:2003:BRB

[457] P. Sprent. Book review: Introduction to Nonparametric Item Response Theory, by K. Sijtsma; I. W. Molenaar. Journal of the Royal Statistical Society. Series D (The Statistician), 52(2): 249–250, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/4128270.

Kemp:2003:BRBn

[458] A. W. Kemp. Book review: Statistical Inference in Science, by D. A. Sprott. Journal of the Royal Statistical Society. Series D (The Statistician), 52(2): 250–251, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/4128271.

Haigh:2003:BRBb

[459] John Haigh. Book review: Biometrika:

One Hundred Years, by D. M. Titterington; D. R. Cox. Journal of the Royal Statistical Society. Series D (The Statistician), 52(2):251–252, ???? 2003. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/4128272.

Eddings:2003:BRB

[460] Wesley D. Eddings. Book review: Conquering Statistics: Numbers without the Crunch, by J. H. Weaver. Journal of the Royal Statistical Society. Series D (The Statistician), 52(2):252–253, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/4128273.

Coolen:2003:BRBb

[461] F. P. A. Coolen. Book review: Elementare Grundbegriffe einer Allgemeineren Wahrscheinlichkeitsrechnung, vol. I, Intervallwahrscheinlichkeit als Umfassendes Konzept, by K. Weichselberger. Journal of the Royal Statistical Society. Series D (The Statistician), 52 (2):253–254, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/4128274.

Bashir:2003:BRB

[462] Saghir Bashir. Book review: First Steps in Statistics, by D. B. Wright. Journal of the Royal Statistical Society. Series D (The Statistician), 52(2): 254–255, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/4128275.

Anonymous:2003:FMb

[463] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 52(2):??, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/4128240.

Smith:2003:DBS

[464] Paul Smith, Mark Pont, and Tim Jones. Developments in business survey methodology in the Office for National Statistics, 1994–2000. Journal of the Royal Statistical Society. Series D (The Statistician), 52(3):257–295, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/4128202.

Senn:2003:FBP

[465] Stephen Senn. Foreword: A blue plaque for Fisher. Journal of the Royal Statistical Society. Series D (The Statistician), 52(3):297–298, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/4128203.

Green:2003:NLW

[466] Peter J. Green. Notes on the life and work of R. A. Fisher. Journal of the Royal Statistical Society. Series D (The Statistician), 52(3):299–301, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/4128204.

Healy:2003:RFS

[467] M. J. R. Healy. R. A. Fisher the statistician. Journal of the Royal Statistical Society. Series D (The Statistician), 52 (3):303–310, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/4128205.

Edwards:2003:RFT

[468] A. W. F. Edwards. R. A. Fisher: Twice Professor of Genetics: London and Cambridge, or ‘a fairly well-known geneticist’. Journal of the Royal Statistical Society. Series D (The Statistician), 52 (3):311–318, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/4128206.

Grafen:2003:FEB

[469] Alan Grafen. Fisher the evolutionary biologist. Journal of the Royal Statistical Society. Series D (The Statistician), 52 (3):319–329, ???? 2003. CODEN ????

ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/4128207.

Yu:2003:QRA

[470] Keming Yu, Zudi Lu, and Julian Stander. Quantile regression: Applications and current research areas. Journal of the Royal Statistical Society. Series D (The Statistician), 52(3): 331–350, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/4128208.

Raats:2003:DCA

[471] V. M. Raats and J. J. A. Moors. Double-checking auditors: A Bayesian approach. Journal of the Royal Statistical Society. Series D (The Statistician), 52(3):351–365, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/4128209.

Hassler:2003:NBC

[472] Uwe Hassler and Thorsten Thadewald. Nonsensical and biased correlation due to pooling heterogeneous samples. Journal of the Royal Statistical Society. Series D (The Statistician), 52(3):367–379, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
4128210.

Karlis:2003:ASD

[473] Dimitris Karlis and Ioannis Ntzoufras. Analysis of sports data by using bivariate Poisson models. Journal of the Royal Statistical Society. Series D (The Statistician), 52(3):381–393, ???? 2003. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/4128211.

Nevzorov:2003:SCS

[474] V. B. Nevzorov, N. Balakrishnan, and M. Ahsanullah. Simple characterizations of Student’s t_{2}-distribution. Journal of the Royal Statistical Society. Series D (The Statistician), 52(3):395–400, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
4128212.

Hibbert:2003:OMJ

[475] Jack Hibbert. Obituary: Michael John Mackey Erritt, 1931–2002. Journal of the Royal Statistical Society. Series D (The Statistician), 52(3):401–402, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
4128213.

Anonymous:2003:BMc

[476] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 52(3):403, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/4128239.

Kemp:2003:BRBo

[477] Freda Kemp. Book review: Sample Survey Principles and Methods, by V. Barnett. Journal of the Royal Statistical Society. Series D (The Statistician), 52(3):404, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/4128214.

Stoyanov:2003:BRBb

[478] Jordan Stoyanov. Book review: Fundamentals of Queueing Networks: Performance, Asymptotics and Optimization, by H. Chen; D. D. Yao. Journal of the Royal Statistical Society. Series D (The Statistician), 52(3):404, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/4128215.

Seuc:2003:BRB

[479] Armando H. Seuc. Book review: Principles and Methods of Social Research, by W. D. Crano; M. B. Brewer. Journal of the Royal Statistical Society. Series D (The Statistician), 52(3):404–405, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
4128216.

Stoyanov:2003:BRBc

[480] Jordan Stoyanov. Book review: From Elementary Probability to Stochastic Differential Equations with MAPLE(R), by S. Cyganowski; P. Kloeden; J. Ombach. Journal of the Royal Statistical Society. Series D (The Statistician), 52 (3):405–406, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/4128217.

Kemp:2003:BRBp

[481] Freda Kemp. Book review: Pascal’s Arithmetical Triangle: The Story of a Mathematical Idea, by A. W. F. Edwards. Journal of the Royal Statistical Society. Series D (The Statistician), 52(3):406, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/4128218.

Best:2003:BRBb

[482] Mark A. Best. Book review: Data Monitoring Committees in Clinical Trials, by S. S. Ellenberg; T. R. Fleming; D. L. DeMets. Journal of the Royal Statistical Society. Series D (The Statistician), 52(3):406–407, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/4128219.

Arnold:2003:BRB

[483] Gillian M. Arnold. Book review: Cluster Analysis, by B. S. Everitt; S. Landau; M. Leese. Journal of the Royal Statistical Society. Series D (The Statistician), 52(3):407–408, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/4128220.

Kemp:2003:BRBq

[484] David Kemp. Book review: Discrete-Event Simulation: Modeling, Programming, and Analysis, by G. S. Flshman. Journal of the Royal Statistical Society. Series D (The Statistician), 52(3): 408–409, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/4128221.

Sisson:2003:BRBb

[485] Scott Sisson. Book review: Modern Statistics for the Life Sciences, by A. Grafen; R. Hails. Journal of the Royal Statistical Society. Series D (The Statistician), 52(3):409–410, ???? 2003. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/4128222.

Swires-Hennessy:2003:BRBc

[486] Ed Swires-Hennessy. Book review: Survey Non-Response, by R. M. Groves; D. A. Dillman; J. L. Eltinge; R. J. A. Little. Journal of the Royal Statistical Society. Series D (The Statistician), 52 (3):410–411, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/4128223.

Imhof:2003:BRBa

[487] Lorens A. Imhof. Book review: Matrix Algebra: Exercises and Solutions, by D. A. Harville. Journal of the Royal Statistical Society. Series D (The Statistician), 52(3):411–412, ???? 2003. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/4128224.

Kemp:2003:BRBr

[488] Freda Kemp. Book review: Independent Component Analysis, by A. Hyvärinen; J. Karhunen; E. Oja; Independent Component Analysis: Principles and Practice, by S. Roberts; R. Everson. Journal of the Royal Statistical Society. Series D (The Statistician), 52(3):412, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/4128225.

Penenberg:2003:BRB

[489] Darryl Neil Penenberg. Book review: Medical Biostatistics, by A. Indrayan; S. Sarmukaddam. Journal of the Royal Statistical Society. Series D (The Statistician), 52(3):412–413, ???? 2003. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/4128226.

Sisson:2003:BRBc

[490] Scott Sisson. Book review: The Basics of S-PLUS, by A. Krause; M. Olsen. Journal of the Royal Statistical Society. Series D (The Statistician), 52(3): 413–414, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/4128227.

Kemp:2003:BRBs

[491] Freda Kemp. Book review: Mathematical and Statistical Methods for Genetic Analysis, by K. Lange. Journal of the Royal Statistical Society. Series D (The Statistician), 52(3):414–415, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/4128228.

Sofroniou:2003:BRB

[492] Nick Sofroniou. Book review: Statistical Concepts: A Second Course for Education and the Behavioral Sciences, by R. G. Lomax. Journal of the Royal Statistical Society. Series D (The Statistician), 52(3):415, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/4128229.

DSouza:2003:BRBb

[493] J. C. D’Souza. Book review: Spline Regression Models, by L. C. Marsh; D. R. Cormier. Journal of the Royal Statistical Society. Series D (The Statistician), 52(3):416, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/4128230.

Larsen:2003:BRBb

[494] Pia Veldt Larsen. Book review: In All Likelihood: Statistical Modelling and Inference Using Likelihood, by Y. Pawitan. Journal of the Royal Statistical Society. Series D (The Statistician), 52 (3):416–417, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/4128231.

Moreno:2003:BRB

[495] Victor Moreno. Book review: Understanding Statistical Concepts Using S-Plus, by R. E. Schumacker; A. Akers. Journal of the Royal Statistical Society. Series D (The Statistician), 52 (3):417, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/4128232.

Rigby:2003:BRBb

[496] Alan S. Rigby. Book review: Crossover Trials in Clinical Research, by S. Senn. Journal of the Royal Statistical Society. Series D (The Statistician), 52 (3):417–418, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/4128233.

Mitchell:2003:BRBc

[497] Lada Mitchell. Book review: Applied Multivariate Statistics for the Social Sciences, by J. Stevens. Journal of the Royal Statistical Society. Series D (The Statistician), 52(3):418–420, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/4128234.

Kemp:2003:BRBt

[498] Freda Kemp. Book review: Statistics at Square One, by T. D. V. Swinscow; M. J. Campbell. Journal of the Royal Statistical Society. Series D (The Statistician), 52(3):420, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/4128235.

Sapatinas:2003:BRB

[499] Theofanis Sapatinas. Book review:

Asymptotic Theory of Statistical Inference for Time Series, by M. Taniguchi; Y. Kakizawa. Journal of the Royal Statistical Society. Series D (The Statistician), 52(3):420–421, ???? 2003. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/4128236.

Woodward:2003:BRB

[500] P. W. Woodward. Book review: Box on Quality and Discovery: With Design, Control and Robustness, by G. C. Tiao; S. Bisgaard; W. J. Hill; D. Peña; S. M. Stigler. Journal of the Royal Statistical Society. Series D (The Statistician), 52 (3):421–422, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/4128237.

Tusell:2003:BRB

[501] Fernando Tusell. Book review: Simplicity, Inference and Modelling, by A. Zellner; H. A. Keuzenkamp; M. McAleer. Journal of the Royal Statistical Society. Series D (The Statistician), 52 (3):422, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/4128238.

Anonymous:2003:FMc

[502] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 52(3):??, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/4128201.

Green:2003:DGS

[503] Peter J. Green. Diversities of gifts, but the same spirit. Journal of the Royal Statistical Society. Series D (The Statistician), 52(4):423–438, ???? 2003. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/4128124.

Holmes:2003:YST

[504] Peter Holmes. 50 years of statistics teaching in English schools: Some milestones. Journal of the Royal Statistical Society. Series D (The Statistician), 52 (4):439–474, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/4128125.

Walker:2003:HMS

[505] Stephen G. Walker. How many samples?: A Bayesian nonparametric approach. Journal of the Royal Statistical Society. Series D (The Statistician), 52(4):475–482, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/4128126.

Papathomas:2003:BUB

[506] Michail Papathomas and Roger J. Hocking. Bayesian updating for binary variables: An application in the UK water industry. Journal of the Royal Statistical Society. Series D (The Statistician), 52(4):483–499, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/4128127.

Luceno:2003:DBA

[507] Alberto Luceño. Dead band adjustment charts with asymmetric off-target costs, deterministic process drift and delayed dynamics. Journal of the Royal Statistical Society. Series D (The Statistician), 52(4):501–514, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/4128128.

Goel:2003:WOE

[508] Prem K. Goel and Josep Ginebra. When is one experiment ‘always better than’ another? Journal of the Royal Statistical Society. Series D (The Statistician), 52(4):515–537, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/4128129.

Brooks:2003:SSD

[509] R. J. Brooks, A. M. Cottenden, and M. J. Fader. Sample sizes for designed studies with correlated binary data. Journal of the Royal Statistical Society. Series D (The Statistician), 52 (4):539–551, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/4128130.

Farrow:2003:PBS

[510] Malcolm Farrow. Practical building of subjective covariance structures for large complicated systems. Journal of the Royal Statistical Society. Series D (The Statistician), 52(4):553–573, ???? 2003.

CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/4128131.

Piepho:2003:FET

[511] Hans-Peter Piepho. The folded exponential transformation for proportions. Journal of the Royal Statistical Society. Series D (The Statistician), 52(4): 575–589, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/4128132.

Hirotsu:2003:ECT

[512] Nobuyoshi Hirotsu and Mike Wright. An evaluation of characteristics of teams in Association Football by using a Markov process model. Journal of the Royal Statistical Society. Series D (The Statistician), 52(4):591–602, ???? 2003. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/4128133.

Anonymous:2003:RCS

[513] Anonymous. Report of the Council for the session 2002–2003. Journal of the Royal Statistical Society. Series D (The Statistician), 52(4):603–678, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/4128134.

Cox:2003:FA

[514] D. R. Cox. Frank anscombe. Journal of the Royal Statistical Society. Series D (The Statistician), 52(4):679–680, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        4128135.

Haberman:2003:PBB

[515] Steven Haberman. Professor Bernard Benjamin, 1910–2002. Journal of the Royal Statistical Society. Series D (The Statistician), 52(4):680–682, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/4128136.

Cowell:2003:PDG

[516] F. A. Cowell. Professor David Gawen Chapernowne. Journal of the Royal Statistical Society. Series D (The Statistician), 52(4):682–683, ???? 2003. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/4128137.

Glasbey:2003:RK

[517] Chris Glasbey, David Elston, and Mike Talbot. Rob kempton. Journal of the Royal Statistical Society. Series D (The Statistician), 52(4):683–684, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/4128138.

Oliver:2003:WFD

[518] Frank Oliver. Wynne Frederick (Derek)

Maunder, 1920–2003. Journal of the Royal Statistical Society. Series D (The Statistician), 52(4):685, ???? 2003. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/4128139.

Burkhardt:2003:DW

[519] Hugh Burkhardt. David Wishart. Journal of the Royal Statistical Society. Series D (The Statistician), 52(4):685–687, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
4128140.

Anonymous:2003:BMd

[520] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 52(4):689, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/4128163.

Hewson:2003:BRBb

[521] Paul Hewson. Book review: Temporal GIS: Advanced Functions for Field-Based Applications, by G. Christakos; P. Bogaert; M. Serre. Journal of the Royal Statistical Society. Series D (The Statistician), 52(4):690, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/4128141.

Rigby:2003:BRBc

[522] Alan S. Rigby. Book review: Statistics in Clinical Practice, by D. Coggan. Journal of the Royal Statistical Society. Series D (The Statistician), 52 (4):690–691, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/4128142.

Kemp:2003:BRBu

[523] Freda Kemp. Book review: Applied Multiple Regression/Correlation Analysis for the Behavioral Sciences, by J. Cohen; P. Cohen; S. G. West; L. S. Aiken. Journal of the Royal Statistical Society. Series D (The Statistician), 52(4):691, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/4128143.

Huson:2003:BRBc

[524] Les Huson. Book review: Statistical Methods for the Analysis of Repeated Measurements, by C. S. Davis. Journal of the Royal Statistical Society. Series D (The Statistician), 52(4): 691–692, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/4128144.

Walters:2003:BRB

[525] Stephen Walters. Book review: Analysis of Longitudinal Data, by P. J. Diggle; P. Heagerty; K.-Y. Liang; S. L. Zeger. Journal of the Royal Statistical Society. Series D (The Statistician), 52 (4):692–694, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/4128145.

Kemp:2003:BRBv

[526] Freda Kemp. Book review: An Introduction to Generalized Linear Models, by A. Dobson. Journal of the Royal Statistical Society. Series D (The Statistician), 52(4):694, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/4128146.

Kemp:2003:BRBw

[527] Freda Kemp. Book review: An Introduction to Sequential Monte Carlo Methods, by A. Doucet; N. de Freitas; N. Gordon. Journal of the Royal Statistical Society. Series D (The Statistician), 52(4):694–695, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/4128147.

Williams:2003:BRB

[528] Sian Williams and Daniel Wright. Book review: Statistics for Psychologists: An Intermediate Course, by B. S. Everitt. Journal of the Royal Statistical Society. Series D (The Statistician), 52(4): 695–696, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/4128148.

Gray:2003:BRB

[529] Alison Gray. Book review: Applied Multivariate Analysis, by B. S. Everitt; G. Dunn. Journal of the Royal Statistical Society. Series D (The Statistician), 52(4):696, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/4128149.

Janacek:2003:BRBa

[530] G. Janacek. Book review: Non-Linear Time Series Models in Empirical Finance, by P. H. Franses; D. van Dijk. Journal of the Royal Statistical Society. Series D (The Statistician), 52(4): 696–697, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/4128150.

Kemp:2003:BRBx

[531] Freda Kemp. Book review: Scan Statistics, by J. Glaz; J. Naus; S. Wallenstein. Journal of the Royal Statistical Society. Series D (The Statistician), 52 (4):697–698, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/4128151.

Kemp:2003:BRBy

[532] Freda Kemp. Book review: Statistical Thinking for Managers, by J. A. John; D. Whitaker; D. G. Johnson. Journal of the Royal Statistical Society. Series D (The Statistician), 52(4):698, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/4128152.

Kemp:2003:BRBz

[533] Freda Kemp. Book review: The Laplace Distribution and Generalizations: A Revisit with Applications to Communications, Economics, Engineering, and Finance, by S. Kotz; T. J. Kozubowski; K. Podgórski. Journal of the Royal Statistical Society. Series D (The Statistician), 52(4):698–699, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/4128153.

Brix:2003:BRBb

[534] Anders Brix. Book review: Geostatistical Simulation, by C. Lantuejoul. Journal of the Royal Statistical Society. Series D (The Statistician), 52(4): 699–700, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/4128154.

Kemp:2003:BRBaa

[535] Freda Kemp. Book review: Thiele:

Pioneer in Statistics, by S. L. Lauritzen. Journal of the Royal Statistical Society. Series D (The Statistician), 52(4):700, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/4128155.

Best:2003:BRBc

[536] Mark A. Best. Book review: Nonlinear Models in Medical Statistics, by J. K. Lindsey. Journal of the Royal Statistical Society. Series D (The Statistician), 52(4):701, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884

                                        (electronic). URL http://www.jstor.
                                        org/stable/4128156.

Barber:2003:BRB

[537] Stuart Barber. Book review: A Handbook of Time-Series Analysis, Signal Processing and Dynamics, by D. S. G. Pollock. Journal of the Royal Statistical Society. Series D (The Statistician), 52(4):701–702, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/4128157.

Starkings:2003:BRB

[538] Susan Starkings. Book review: Basic Statistical Methods and Models for the Sciences, by J. Rosenblatt. Journal of the Royal Statistical Society. Series D (The Statistician), 52(4):702–703, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        4128158.

Janacek:2003:BRBb

[539] G. Janacek. Book review: Testing for Normality, by H. C. Thode. Journal of the Royal Statistical Society. Series D (The Statistician), 52(4):703, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/4128159.

Imhof:2003:BRBb

[540] Lorens A. Imhof. Book review: Matrix Calculus and Zero–One Matrices: Statistical and Econometric Applications, by D. A. Turkington. Journal of the Royal Statistical Society. Series D (The Statistician), 52(4):703–704, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/4128160.

Kemp:2003:BRBab

[541] Freda Kemp. Book review: Modern Applied Statistics with S, by B. Venables; B. Ripley. Journal of the Royal Statistical Society. Series D (The Statistician), 52(4):704–705, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/4128161.

Tolvi:2003:BRBc

[542] Jussi Tolvi. Book review: Modeling Financial Time Series with S-Plus, by E. Zivot; J. Wang. Journal of the Royal Statistical Society. Series D (The Statistician), 52(4):705, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/4128162.

Anonymous:2003:VI

[543] Anonymous. Volume information. Journal of the Royal Statistical Society. Series D (The Statistician), 52(4):706–712, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
4128122.

Anonymous:2003:FMd

[544] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 52(4):??, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/4128123.